### v1

# bibliography.bib

[…]
bibliography.bib
3DES
Triple DES
AES
Advanced Encryption Standard
AMD
Advanced Micro Devices
ANF
Algebraic Normal Form
AoT
Ahead of Time
API
Application Programming Interface
ARM
Advanced RISC Machines
ARX
Addition, Rotate, XOR
ASK
Amplitude-Shift Keying
ASIC
Application Specific Integrated Circuit
BL
Bootloader Enable
BOR
Brown-Out Reset
BPSK
Binary Phase Shift Keying
CBC
Cipher Block Chaining
CBS
Critical Bootloader Section
CGM
Continuous Glucose Monitoring System
CMOS
Complementary Metal Oxide Semiconductor
CNN
Convolutional Neural Network
COPACOBANA
Cost-Optimized Parallel Code Breaker and Analyzer
CPA
Correlation Power Analysis
CPU
Central Processing Unit
CTR
Counter (mode of operation)
CRC
Cyclic Redundancy Check
CRP
Code Readout Protection
DES
Data Encryption Standard
DC
Direct Current
DDS
Digital Direct Synthesis
DFT
Discrete Fourier Transform
DMA
Direct Memory Access
DNNL
Deep Neural Network Library
DoS
Denial-of-Service
DFA
Differential Fault Analysis
DPA
Differential Power Analysis
DRAM
Dynamic Random-Access Memory
DRM
Digital Rights Management
DSO
Digital Storage Oscilloscope
DSP
Digital Signal Processing
DST
Digital Signature Transponder
DUT
Device Under Test
ECB
Electronic Code Book
ECC
Elliptic Curve Cryptography
ECU
Electronic Control Unit
EDE
Encrypt-Decrypt-Encrypt (mode of operation)
EEPROM
Electrically Erasable Programmable Read-Only Memory
EM
electro-magnetic
FFT
Fast Fourier Transform
FCC
Federal Communications Commission
FIR
Finite Impulse Response
FIVR
Fully Integrated Voltage Regulator
FNN
Feedforward Neural Network
FPGA
Field Programmable Gate Array
FSK
Frequency Shift Keying
GMSK
Gaussian Minimum Shift Keying
GPIO
General Purpose I/O
GPU
Graphics Processing Unit
HD
Hamming Distance
HDL
Hardware Description Language
HF
High Frequency
HMAC
Hash-based Message Authentication Code
HW
Hamming Weight
IC
Integrated Circuit
ID
Identifier
ISM
Industrial, Scientific, and Medical (frequencies)
IIR
Infinite Impulse Response
IP
Intellectual Property
IoT
Internet of Things
IV
Initialization Vector
JTAG
Joint Test Action Group
LF
Low Frequency
LFSR
Linear Feedback Shift Register
LLM
Large Language Model
LQI
Link Quality Indicator
LSB
Least Significant Bit
LSByte
Least Significant Byte
LUT
Look-Up Table
MAC
Message Authentication Code
MF
Medium Frequency
MMU
Memory Management Unit
MITM
Man-In-The-Middle
MSB
Most Significant Bit
MSByte
Most Significant Byte
MSK
Minimum Shift Keying
MSR
Model Specific Register
NLFSR
Non-Linear Feedback Shift Register
NLF
Non-Linear Function
NFC
Near Field Communication
NRZ
Non-Return-to-Zero (encoding)
NVM
Non-Volatile Memory
OOK
On-Off-Keying
OP
Operational Amplifier
OTP
One-Time Password
PC
Personal Computer
PCB
Printed Circuit Board
PhD
Patiently hoping for a Degree
PKE
Passive Keyless Entry
PKES
Passive Keyless Entry and Start
PKI
Public Key Infrastructure
PMBus
Power Management Bus
PoC
Proof-of-Concept
POR
Power-On Reset
PPC
Pulse Pause Coding
PRNG
Pseudo-Random Number Generator
PSK
Phase Shift Keying
PWM
Pulse Width Modulation
RAPL
Running Average Power Limit
RDP
Read-out Protection
RF
Radio Frequency
RFID
Radio Frequency IDentification
RKE
Remote Keyless Entry
RNG
Random Number Generator
ROM
Read Only Memory
ROP
Return-Oriented Programming
RSA
Rivest Shamir and Adleman
SEV
Secure Encrypted Virtualization
SCA
Side-Channel Analysis
SDR
Software-Defined Radio
SGX
Software Guard Extensions
SNR
Signal to Noise Ratio
SHA
Secure Hash Algorithm
SHA-1
Secure Hash Algorithm 1
SHA-256
Secure Hash Algorithm 2 (256-bit version)
SMA
SubMiniature version A (connector)
SMBus
System Management Bus
I2C
Inter-Integrated Circuit
SPA
Simple Power Analysis
SPI
Serial Peripheral Interface
SPOF
Single Point of Failure
SoC
System on Chip
SVID
Serial Voltage Identification
SWD
Serial Wire Debug
TCB
Trusted Computing Base
TDX
Trust Domain Extensions
TLB
Translation Lookaside Buffer
TEE
Trusted Execution Environment
TMTO
Time-Memory Tradeoff
TMDTO
Time-Memory-Data Tradeoff
TrustZone
TrustZone
RSSI
Received Signal Strength Indicator
SHF
Superhigh Frequency
UART
Universal Asynchronous Receiver Transmitter
UHF
Ultra High Frequency
UID
Unique Identifier
USRP
Universal Software Radio Peripheral
USRP2
Universal Software Radio Peripheral (version 2)
USB
Universal Serial Bus
VHF
Very High Frequency
VLF
Very Low Frequency
VHDL
VHSIC (Very High Speed Integrated Circuit) Hardware Description Language
VM
Virtual Machine
VR
Voltage Regulator
WLAN
Wireless Local Area Network
XOR
Exclusive OR
IR
Intermediate Representation
OCD
On-Chip Debug
OS
Operating System
VM
Virtual Machine
ML
Machine Learning
DVFS
Dynamic Voltage and Frequency Scaling
MMIO
Memory-Mapped I/O
Activation Functions Considered Harmful: Recovering Neural Network Weights through Controlled Channels
Jesse Spielman and David Oswald and Mark Ryan and Jo Van Bulck School of Computer Science
University of Birmingham
Birmingham, UK
jxs1366@student.bham.ac.uk
School of Computer Science
University of Birmingham
Birmingham, UK
d.f.oswald@bham.ac.uk
School of Computer Science
University of Birmingham
Birmingham, UK
m.d.ryan@bham.ac.uk
DistriNet
KU Leuven, Belgium
Leuven, Belgium
jo.vanbulck@cs.kuleuven.be
Abstract
With high-stakes machine learning applications increasingly moving to untrusted end-user or cloud environments, safeguarding pre-trained model parameters becomes essential for protecting intellectual property and user privacy. Recent advancements in hardware-isolated enclaves, notably Intel SGX, hold the promise to secure the internal state of machine learning applications even against compromised operating systems. However, we show that privileged software adversaries can exploit input-dependent memory access patterns in common neural network activation functions to extract secret weights and biases from an SGX enclave. Our attack leverages the SGX-Step framework to obtain a noise-free, instruction-granular page-access trace. In a case study of an 11-input regression network using the Tensorflow Microlite library, we demonstrate complete recovery of all first-layer weights and biases, as well as partial recovery of parameters from deeper layers under specific conditions. Our novel attack technique requires only 20 queries per input per weight to obtain all first-layer weights and biases with an average absolute error of less than 1%, improving over prior model stealing attacks.
Additionally, a broader ecosystem analysis reveals the widespread use of activation functions with input-dependent memory access patterns in popular machine learning frameworks (either directly or via underlying math libraries). Our findings highlight the limitations of deploying confidential models in SGX enclaves and emphasise the need for stricter side-channel validation of machine learning implementations, akin to the vetting efforts applied to secure cryptographic libraries.
Index Terms:
Machine Learning Security, Side Channel Attacks, Trusted Execution, Confidential Computing, SGX
1 Introduction
Training and deploying Machine Learning (ML) models, in particular neural networks, often requires a massive investment in computational resources. Hence, the threat of model stealing attacks is significant, because they allow an attacker to abscond with an ML application, causing substantial financial damage as well as security and privacy risks. Attacks could be used for industrial espionage (in the case of a network that solves a difficult problem, e.g., a sophisticated automated trading system), or to ease the creation of adversarial examples to defeat a classification system, e.g., for spam filtering or intrusion detection. An application that incurs a fee to access per interaction could also be duplicated to remove this restriction. While existing model stealing attacks show how to duplicate neural networks essentially via brute force, they scale poorly for increasingly complex networks and require detailed access to output probabilities or physical access for measurements such as power consumption.
Because ML (both training and inference) requires substantial computation, there is a strong focus on performance. Further, because of the hardware requirements of ML training and inference (such as in the case of an AI model which backs an end user-accessible web interface), there is an increasing need to host ML workloads in the cloud, e.g., for scaling or to use the latest Central Processing Unit (CPU) and Graphics Processing Unit (GPU) hardware. However, this move off-premises and onto (often shared) hardware introduces a number of security concerns. To mitigate the risk, substantial work has been done using Trusted Execution Environments such as Intel Software Guard Extensions (SGX) or AMD Secure Encrypted Virtualization (SEV) to protect ML workloads running on third-party hardware [russinovich2023confidential, intel_secure_ml, kumar23, li2024coreguard, kethireddy2024secure].
A core component of neural networks are activation functions, chosen by developers (alongside other architectural decisions such as the number of layers or the learning rate) to introduce non-linearity. These functions are generally computed once for each neuron based on the sum of the multiplication of all incoming inputs by their weights (plus a bias term). Often, these functions are provided by a ML framework such as Tensorflow or PyTorch, which all include routines for training, running inference, and deploying models. It is important to note that these framework-provided activation functions usually rely on basic mathematical functions such as std::max() or exp(), often provided by the system’s standard libraries (e.g., glibc or musl) and optimised for speed, not side-channel resistance. TEEs like Intel SGX often supply their own (restricted) standard library (e.g., SGX’s tlibc), which provides implementations known to work within the special enclaved environment. In summary, ML frameworks inherit their security properties from the used standard library math functions.
Research into side-channel attacks (both with physical access as well as remotely exploitable side channels like timing) has led to the “hardening” of certain security sensitive algorithms, especially in cryptographic libraries. However, less attention has been paid to the side-channel security of ML libraries, even though their parameters (weights and biases) are today often akin to cryptographic secrets in sensitivity. Especially in the context of TEE-protected ML workloads, the strengthened side-channel adversary model has not been sufficiently studied.
In this paper, we show how pervasive data-dependent memory access patterns in activation functions across many ML frameworks can lead to the deterministic recovery of hidden model parameters when deployed in a TEE. Particularly, we develop a novel methodology to recover partial weights and biases from the first (and deeper) layer(s) of a victim network when provided with an instruction-granular page-access trace. As a practical case-study attack, we deploy a Tensorflow Microlite model inside an SGX enclave and extract deterministic, instruction-granular page access traces using the SGX-Step [van_bulck_sgx-step_2017] framework. We demonstrate successful recovery of each weight and bias of the network to more than 5 decimal places accuracy with 55 invocations of the network, or with less than 1% average error with around 20 invocations after an initial calibration phase. Our end-to-end attack outperforms Tramèr-style model stealing attacks [tramer_stealing_2016], which require around 100 invocations for each parameter in simpler classification networks. Considering the wider ecosystem, we survey widely used ML libraries and find widespread use of secret-dependent access patterns in activation functions. Furthermore, we discuss to what extent our attack vector is applicable to other popular TEE architectures.
Contributions
Summarised, our contributions are:
  • A novel methodology to recover weights and biases from partial memory-access side-channel traces.
  • An end-to-end case study of practical memory-trace extraction and accurate weight/bias recovery from Tensorflow Microlite running inside an SGX enclave.
  • A comprehensive survey of input-dependent accesses in common activation functions across popular ML and standard libraries.
  • An open-source Tensorflow Microlite SGX benchmark for future work on attacks and mitigations.
Ethical Considerations
This work contributes to better understanding limitations of widely used TEE technology. Our attacks exploit the well-documented issue of precise monitoring of enclave memory accesses using tools like SGX-Step. All experiments were conducted on proof-of-concept implementations on our own local machines.
2 Background and Related Work
Trusted Execution Environments
One of most commonly used and widely studied TEEs is Intel’s SGX, which, even though discontinued for client CPUs, is widely supported on Xeon Scalable server CPUs and now marketed for use cases such as secure ML deployments [intel_secure_ml]. SGX-secured code is executed in a hardware-protected ‘enclave’ (with its memory encrypted) such that even the root user cannot gain access. SGX does not make any guarantees about protections against timing and memory-based side channels [noauthor_intel_nodate], and hence is vulnerable to cache and page fault attacks [nilsson_survey_2020]. SGX also allows secrets to be securely loaded into the enclave remotely, and subsequently be (un)sealed for local storage. This allows data of the enclave (such as a pre-trained ML model) to be also protected while at rest.
Other TEEs of note include ARM’s TrustZone which brings the concept of TEE to e.g., mobile devices for isolating certain software from the user, for example, in Digital Rights Management (DRM) applications [arm_tz]. Intel’s Trust Domain Extensions (TDX[intel_tdx] and AMD’s SEV [amd_sev] are TEEs that build on application-level TEEs which provide hardware-backed isolation guarantees at the level of entire Virtual Machines. Their approach is intended to make it easier for developers to take advantage of TEE security features, because applications do not need to be ported or modified for the enclaved security model.
Though not a vulnerability per se, a consequence of SGX’s adversary model is that it allows an attacker to have precise control over the enclave’s execution. This gives rise to tools like SGX-Step [van_bulck_sgx-step_2017], which allows an attacker to perfectly single-step enclaved code by causing the enclave to be interrupted, suspended, and then resumed rapidly. When paired with a more conventional side channel that can probe memory pages to detect other accesses, a precise record of a program’s execution (or trace) with respect to those pages can be recorded.
Model Stealing Attacks
[…]
Side-Channel Attacks on ML Implementations
[…]
In contrast, side-channel attacks that recover the full model with physical access to the hardware and measurements of electro-magnetic (EM) emanation or power consumption have been further developed [Hua18, Batina19, Yu20, horvath2024barracuda]. However, as these attacks require measurements on or in close proximity to the targeted chips, they are of less relevance to datacenter applications, where physical access to the underlying servers is typically closely guarded.
Hence, in this paper, we focus on software-based side-channel attacks to recover weights/biases without physical access and without access to the output probabilities.
Controlled Channel Attacks
For SGX, Intel explicitly places the burden of preventing secret-dependent memory accesses for security-critical code on the developer [Wenhao17]. In practice, this means that enclave developers should avoid data-dependent patterns of the form array[x] or if(x) func(), where x depends on a secret. It is important to note that in practice, such secret-dependent access patterns may only become evident at the machine code level, as optimising compilers might e.g., opt to compile “branchless” code into code with conditional branches, and vice versa.
[…]
Figure 1: Victim enclaves exhibiting tight memory-access patterns can be precisely interrupted at instruction-level granularity using SGX-Step allowing to extract deterministic page-access count traces.
The open-source SGX-Step framework [van_bulck_sgx-step_2017, vanbulck2023sgxstep] allows for precise single-stepping of production enclaves using privileged x86 APIC timer interrupts, such that page access patterns can be deterministically monitored for every enclave instruction, giving precise insights into the operation of a victim enclave. By conservatively under-estimating the APIC timer interval, SGX-Step results in either zero or single-steps, but always avoids multi-steps [vanbulck2023sgxstep, Table 5]. Subsequent works [VanBulck18, moghimi2020copycat, aldaya2020one, van_bulck_tale_2019, constable2023aexnotify] showed that the accessed (A) bit in the enclave code page deterministically distinguishes single-steps (enclave instruction retired; A=1) from zero-steps (A=0). Intel has acknowledged perfect single stepping in a recent paper [constable2023aexnotify], which also explains the underlying mechanism to stall the CPU pipeline using microcode assists. To date, instruction-granular page-access traces extracted with SGX-Step have been repeatedly abused to reliably exploit enclave interface vulnerabilities [van_bulck_tale_2019] or to reconstruct cryptographic key material [moghimi2020copycat, aldaya2020one]. In the following, we show that such traces are sufficient to recover weights from real-world neural network libraries running inside a production SGX enclave.
3 System and Adversary Model
We adhere to Intel’s standard SGX threat model, where the adversary has full root access on the victim machine, as is e.g., the case when deploying an ML model on external cloud infrastructure. Notably, we do not require physical access to the machine for our attack. Privileged software adversaries can use SGX-Step (cf. Section 2) to precisely single-step a target enclave and obtain instruction-granular page access traces. We discuss generalizations to other TEEs in Section 8
We assume the network’s weights and biases are confidential (sealed), while the enclave source code is known. The latter is a common assumption in SGX attacks [xu_controlled-channel_2015, van_bulck_sgx-step_2017, VanBulck17, Brasser17, shinde2016preventing] and the default case in the Intel SGX architecture and SDK. Non-standard confidential-code deployments only add “security through obscurity” by requiring an additional, out-of-scope reversing phase [puddu2024lack, yan_cache_2020].
We assume a pre-trained ML network has been securely loaded into an SGX enclave to perform inference. We do not consider the security of the model training process, which is orthogonal to our attacks, but believe it to be an interesting avenue for future research. Inputs are passed into the enclave and the output is returned outside the enclave through the standard ecall interface. Based on the security guarantees of SGX, an ML developer would assume that their network is protected against duplication, as the inference is handled entirely within an enclave, and weights are never stored outside.
We focus on Feedforward Neural Networks because they effectively illustrate the problem we exploit and are easily deployable in SGX enclaves (e.g., without dynamic linking or requiring libc) using Tensorflow Microlite. We also assume the network architecture (at least the number of neurons per layer and activation functions) is evident from the source code, known (as in the case of a well-known architecture or a transfer-learning model), or can be recovered via a cache attack [yan_cache_2020].
4 Vulnerable Activation Functions
In this paper, we mostly consider activation functions that include a call or calls to a functions in the exp() family, e.g., sigmoid(), tanh(), and expf(). This is based on our initial observation that typical implementations of exp() contain a number of input-dependent branches with different return cases. To give an example, LABEL:lst:expf_in_tlibc shows some different code paths of the implementation of expf() in SGX’s tlibc [^1] https://github.com/intel/linux-sgx/blob/7385e1/sdk/tlibc/math/e_expf.c.
1float expf(float x)
2{
3 // ...
4 if(hx >= 0x42b17218) { /* if x>=88.721... */
5 if(hx>0x7f800000) return x+x; /* NaN */
6 if(hx==0x7f800000) return (xsb==0)? x:0.0;
7 if(x > o_threshold) return ...; // overflow
8 if(x < u_threshold) return ...; // underflow
9 }
10 if(hx > 0x3eb17218) {
11 if(hx < 0x3F851592) {
12 // ...
13 } else {
14 // ...
15 }
16 }
17 else if(hx < 0x31800000) { /* when x<2**-28 */
18 if(huge+x>one) return ...; // inexact
19 }
20 // ...
21}
Listing 1: Excerpt from expf() in tlibc.
Observing the memory access/execution patterns, e.g., with SGX-Step, allows an attacker to learn which return case an input results in, and thus obtain information on the input to exp(). Note that while several variants of exp() implementations exist (both in terms of algorithms for calculation and in terms of single and double precision floats), many of them feature the same structure, where several cases are handled with different return statements.
Attacking an activation function containing a call to exp() is similar to attacking exp() directly. One must be careful to note the sign of the input (as in the case of the $-x$ in sigmoid()’s $\exp\left(-x\right)$) and search in the correct direction. In the case of functions like tanh(), which include multiple discrete calls to exp(), the code to process the memory access traces can not assume a one-to-one relationship between a call to exp() and a single neuron’s activation. Another important consideration is the expressive power of the function, which we discuss in more detail in Section 5.3.
We summarise common activation functions in Table I, and discuss their security in Section 7.2.
Despite our focus on the exp() family of functions, any function exhibiting measurable input-dependent variation in instruction counts (or memory access patterns) is likely vulnerable.
As long as the attacker can force a vulnerable function to process inputs that fall into multiple distinguishable classes, it is possible to search for the boundary points between these cases, as we discuss in detail in Section 5. In particular, relu() does not rely on a call to exp(), but to conditional logic or a call to std::max(). While relu() is ubiquitous, especially in ML-aided computer vision applications, transformers and other modern Large Language Models use a variety of exp()-based activation functions such as softmax(), tanh() and GELU() in addition to relu(). We discuss the vulnerability of relu() in Section 7.2.1.
TABLE I: Survey of Potentially Vulnerable Activation Functions.
Activation Function Output Range Contains exp() Contains max()
sigmoid() $(0,1)$
tanh() $(-1,1)$
softplus() $(0,\infty)$
ELU() $(-\alpha,\infty)$
SELU() $(-\lambda*\alpha,\infty)$
relu() $(0,\infty)$
4.1 ML Framework Functions
Thus far, our discussion of the functions used for activation functions have focused on low level mathematical used by ML framework developers inside of higher level framework code. A Tensorflow neuron with tanh() activation might not, for instance, directly return the output of a call to math library’s implementation of tanh(); there might be wrapping or processing of the result. This is of interest as we can also exploit input-dependent behavior at this higher framework or application level.
An example of this is the code to Tensorflow lite’s sigmoid() function (called Logistic() in the source), shown in LABEL:lst:sigmoid_in_tflite in Appendix B. We can see that on top of any side-channel leaks due to calls to std:exp(), a higher-level set of memory access patterns and instruction count cases are introduced by a conditional routing to different functions (or simply a return of 1) based on the input. In this way, this sigmoid() function can actually be attacked at two levels which we will describe in Section 6.2.
5 Methodology
First, we describe the overall concept for our attack (Section 5.1) against a neural network. We then explain in more detail how to attack the first layer (Section 5.2) and finally the subsequent layers (Section 5.3).
5.1 Attack Concept
The starting point for our attack is the above observation (cf. Section 4) that, unlike in cryptographic libraries, the math functions that underpin many ML activation functions have not been hardened to ensure they execute in constant time and without input-dependent branches and memory access patterns [^2] A classic example in the cache timing literature is Look-Up Tables that may be accessed in deterministic patterns based on the function inputs.. We believe this work has not been undertaken in the context of ML due to performance concerns or because security is seen as less of a focus.
By examining the underlying functions statically, dynamically, or by reading the code (if available), we can deduce the possible branches through the function and the timing as well as memory access patterns therein. Also of interest are cases where certain input values are handled differently (via multiple return statements). A special case of those are early returns, where certain inputs cause the function to return before the full output is computed.
$i_{1}$$w_{1}$$i_{2}$$w_{2}$$\Sigma$$f(\Sigma)$ Activation function $output$ Bias $b$
Figure 2: Visualization of a single 2-input neuron. We use $\Sigma$ in this paper to represent the sum of all $i_{i}*w_{i}$ and $b$ passed into an activation function.
Given that such code patterns take different numbers of instructions (and exhibit different memory access patterns) depending on the input, we can more generally assert that there is a correlation between different code paths and the execution trace. Though these differences might be subtle (e.g., a single instruction), they are measurable given a suitable side channel. Thus, if an attacker can determine which branch a function has taken, they learn something about the input or argument to that function.
As mentioned before, one function with these properties is expf(). Figure 3 shows the 11 different return cases of this function based on input and their associated CPU instruction counts (collected using gdb). We note that we found other optimised implementations of expf() that all use a similar algorithm and thus exhibit the same input-dependent memory access patterns. We focus here on the one used by SGX tlibc. For ease of discussion we labeled these cases using the names shown. For example, expf(4) puts us in case Normal (+)(+)( + ), whereas expf(110) results in the Overflow case. We define a threshold (shown in Figure 3 as dotted black lines) as an input value that is on the precise border of two output classes. For example, $88.72168$ is the value between Normal (+)(+)( + ) and Overflow.
We can also use knowledge of these regions and thresholds in the other direction: for example, if we know that an execution of expf() takes 17 CPU instructions (placing us somewhere in the Overflow region), we know the input to that function call must be greater than $88.72168$. We note that the cases for Overflow and Underflow are both early returns and as such their instruction counts are much lower than the other ‘normal’ returns which actually perform some sort of unique computation.
We can use these threshold values to leak information about the precise value of a partially controlled input. If we consider the scenario where an input to expf() is being multiplied by a hidden constant $c$ and we have access to the instruction count of the execution, we can reveal $c$ by finding (via e.g., a binary search) a threshold value between two instruction count classes and then dividing that value by our input. For example, if we find that the input $3.2164$ multiplied by $c$ leads to threshold value $88.72168$, $c$ must be $\nicefrac{{88.72168}}{{3.2164}}\approx 27.5842$. We therefore define $3.2164$ as a convergence point: an input to the system that causes a threshold value and thus a leak of the precise input to the function.
Figure 3: Visualization of expf() CPU instruction count regions by input. The Underflow and Overflow regions have different instruction counts: 18 and 17 respectively. Note also the symmetrical regions of Inner1, inner2, and Inner3.
We now extend the above into a chosen input attack against a toy ML network. Consider an attack against a simple neural network with only a single neuron, as shown in Figure 2. The activation function is simply expf(). The goal is to recover $w_{1}$, $w_{2}$, and $b$. We are free to choose values for $i_{1}$ and $i_{2}$. We first note that the input to a neuron’s activation function is the sum of the products of various user-controlled inputs and their hidden weights, plus a hidden bias term ($\Sigma$ in Figure 2); this is similar to the case described previously except we have two hidden constants multiplied by two different inputs and an additional term, the bias, that is never multiplied with our inputs.
If there was no bias term, we could simply use a variant of the approach described above and iterate through each input (setting all other inputs to zero, allowing us to focus on one weight at a time) and recovering each parameter independently. The addition of the bias term adds another unknown which means we can only find through constraints on the value of the hidden weights and bias:
$i_{1}*w_{1}+0*w_{2}+b=88.72168$
$\Rightarrow w_{1}=\frac{88.72168-b}{i_{1}}$
To solve for all three unknowns, we find enough sets of values (one for each input to the neuron) to generate a system of equations to solve for as many unknowns as we have (number of weights plus bias term), for example:
[…]
$\cancel{0*w_{1}}+i_{2}*w_{2}+b$ $=threshold_{2}$
$i_{3}*w_{1}+i_{4}*w_{2}+b$ $=threshold_{3}$
[…]
Finally, if we replace exp() above with a different function (such as sigmoid()) that also leaks threshold values, the only thing that changes is how to conduct the binary search: the recovery procedure is otherwise identical. Building on this concept, we will show how it is possible to recover the hidden parameters of more complex networks to a high degree of accuracy.
5.2 Full Attack on First Layer
Continuing on from Section 5.1, we discuss how to attack the first layer of a production neural network. As before, the goal of this phase is to find a number of convergence point sets equal to the number of inputs plus one in order to have a system of equations we can solve to recover the weights and bias.
The first layer may be a special case if (without loss of generality) there is no filtering or normalization which pre-processes the input values. If this is the case, we can choose any values we want to feed into the network (we can not immediately inject values deeper into the network). If there is filtering or normalization, however, we treat the first layer like all subsequent layers.
If we can inject arbitrary values into the first layer, we iterate over every neuron in the first layer and recovery its hidden parameters in 2 phases: calibration and binary search.
5.2.1 The Calibration Phase
In the first phase, we quickly produce sensible starting points and context for the second phase of the attack. We prepare a test input to the network that is all small non-zero values but set the first value of the array to a large value. If we run this test input and detect that the neuron overflowed, we know that this input is large enough to cause an overflow and that the weight between the first input and this neuron must be positive. If instead we produce an underflow, we know that the weight must be negative. If we are not in an overflow or underflow state (perhaps owing to a low magnitude weight). In this case, we simply multiply the large value in our test input by 10 and try again. We continue in this way until we find the sign and magnitude required to over/underflow this neuron from each input to the network. Note that we do not actually binary-search towards a threshold between two cases here; we merely attempt to reach an overflow/underflow case. In this case, we do not need to find a threshold for the instruction count leak to be informative.
Recovering the signs for each weight is useful since some of the cases (most of the “normal” ones) are symmetrical about zero, meaning we cannot learn the sign of the input only based on the state. Recovering the sign before we begin the binary search means we always know exactly where we are in the state space. Recovering a value that overflows/underflows the neuron is useful to accelerate the attack. Though we could choose the maximum float value for the starting point of our binary search, the lower the initial guess, the fewer steps the search has to take. Searching "up" in the calibration step and then searching down in the second phase reduces the number of invocations needed.
5.2.2 The Binary Search Phase
Armed with the information collected in the calibration phase, we continue the attack against the same neuron in phase 1. Now, we can binary search to find the desired number of convergence point sets for this neuron. For as many convergence sets as we need, we start by generating an input set that is made of $n$ locked random values (one for each input) except that the $i$’th input is instead set to the large positive value recovered during calibration. We feed the input into the network and note the return case by interpreting the CPU instruction counts for each neuron. Even if some of the weights have different signs, the overall $\Sigma$ should be dominated by the large input multiplied by its weight. If the neuron has overflowed, our input has caused the neuron to enter the state above the threshold between normal (+) and overflow and should search down. If the neuron has instead underflowed, we must be in the boundary between underflow and normal (-). If the input is negated before entering exp() as in sigmoid(), the direction to search should be reversed. We continue the binary search on that $i$’th input (searching back up if we fall below an underflow/overflow case in terms of magnitude) until we hit the desired search depth or our search does not change, because we reached the limit of decimal places for a floating point number. The tradeoff between depth and speed is shown in Table II.
Using this algorithm, we can generate as many novel convergence sets as we require since the unchanging members of the input set are all randomised between inputs. It is worth noting that the system of equations does not produce usable results if the values on the right hand side are all the same threshold, so targeting multiple thresholds is required. One quick way to find more thresholds for the system of equations is to flip the signs of the inputs to ensure both the underflow-to-normal and overflow-to-normal thresholds are included in the solution set. Note however that it is not required to target the thresholds between underflow/normal and normal/overflow; these are just the easiest to find since we can force an under/overflow and then work our way up/down until we find the normal region. There are other thresholds within exp() (as shown in Figure 3) that can be exploited if they can be reached so long as we know the sign of the weight and thus in which precise state we are in.
Once we have the requisite number of equations, we can solve the system of equations as described in Section 5.1 to recover the weights/biases for this first neuron. We then perform the exact same two-part algorithm on each remaining neuron until we solve the entire layer. The output from our recovery tool as it recovers one neuron is depicted in Appendix A.
5.3 Extending the Attack to Deeper Layers
In order to extend the attack into subsequent layers, we must solve two key problems. First, how can we ‘unwrap’ the effect of already-solved early layers to insert arbitrary inputs beyond them, deeper into the network? And second, is it always possible to generate those values?
Unwrapping Solved Layers
Our attack methodology for the first layer relies on being able to precisely craft a set of “target” in order to search towards threshold points. In order to be able to do this for deeper layers, we first have to “unwrap” all the recovered layers between the input layer and the layer we target. In the following set of equations, we show how to rearrange the normal operation of a neuron to isolate the input as opposed to the target.
$act(input*W+b)$ $=target$
$input*W+b$ $=act^{-1}(target)$
$input*W$ $=act^{-1}(target)-b$
[…]
$input$ $=(act^{-1}(target)-b)*W^{-1}$
Note that it is also possible to do this without matrix inversion for greater numerical efficiency, e.g., using numpy.linalg.lstsq(W, act_inv(target)-b).
In these equations, $act$ and $actinv$ are known since we know the architecture of the network and $W$ and $b$ are known because we already solved this layer. Therefore, given a solved layer’s hidden parameters, a desired ‘target’ set of values, and the appropriate inverse activation function, we can calculate the input we must feed into the layer to produce a desired target as the output of that layer. Note also that the above assumes the activation function is invertible without a substantial loss of accuracy. Inverting these functions might lead to an inability to pass certain values forward, which we discuss below.
Expressive Power
Even if we can algebraically calculate how to pass arbitrary values through a solved layer, subsequent layers are harder to attack because we are limited by the “expressive power” (see Table I) of the neurons before the target layer in terms of both magnitude and sign. There are two consequences to this.
The first is some values may be impossible to generate in subsequent layers. For example, since the output of sigmoid() is always positive, baring negative biases, it would be impossible to send negative values into the next layer. Similarly, relu() can never produce a negative output. The second is that certain thresholds such as those between expf()’s overflow/underflow and normal cases have a high magnitude ($\approx 88.723{-}103.972$) relative to generally smaller magnitude weights and biases. sigmoid() has a very small output range ($0{-}1$) making it impossible to produce a large enough $\Sigma$ in the next layer to ‘reach’ one of these two thresholds unless either the network is very dense with small-magnitude positively weighted neurons or has some positive and large magnitude ($>1$) weights connected to it. This suggests denser networks are more expressive and therefore more vulnerable. In both of these cases, we rely on certain key hidden parameters, which means that certain networks might not be vulnerable to our attack. We however note that security guarantees should never assume a certain network architecture.
Attack Outline
Nevertheless, extending the attack into deeper layers is possible with certain caveats. There is no need to focus only on the high magnitude thresholds we previously discussed between the underflow/overflow cases and the normal one; there are six in the range from $\approx$ -1.04–1.04 (see Figure 3). By searching anywhere within this region (e.g., near the subnormal case as in [Gongye20]) we can, as before, build convergence sets for our inputs and solve for the hidden parameters. Because we might pass through a low expressiveness activation function like sigmoid (output range from 0–1), which is likely being further reduced by multiplication with a weight with magnitude <1, we might not cross a single threshold point (the presence of the bias as a component of $\Sigma$ means that we might be shifted away from zero far enough that we cannot reach both sides of the subnormal case).
As such, we instead start with a grid search to scan (at some resolution up to the attacker which trades off speed and the chances of finding a narrow instruction count class) for multiple instruction count classes. As with the binary search, we start by creating an input set that is made up of $n$ locked random values (one for each input) except that the (randomly chosen) $i$’th input is nominated as a dynamic value, initially set to a high magnitude positive number. We scan the dynamic value from its initial value down to a high magnitude negative number, we can check to see if encounter multiple return cases. If we are able to reach multiple classes, we can then perform a binary search between those classes to isolate the threshold and build a convergence point set, as before.
6 Implementation and Orchestration
We trained several neural networks using “full” Tensorflow (in Python). In this section, we will first introduce a proof-of-concept attack on a case-study victim network with an expf() activation function. Next, we show how to attack a more realistic activation function and start to recover partial information from deeper layers.
6.1 Regression Model with expf()
Model
We first focus on a regression model to predict home insurance costs, which was trained on the “Medical Cost Personal Dataset” [choi18]. The network has three hidden layers. The first hidden layer contains 100 neurons that use the Exponential() activation function which is just a direct call to expf(). There are 11 input nodes, meaning that there are in total 1200 parameters (1100 weights and 100 biases) that we target. After the first layer, there is a second hidden layer of 10 neurons which uses a relu() based activation function. A final layer contains a single relu() neuron.
For this first network, we only focus on the first hidden layer. Despite that, we note that:
  1. 1.
    These choices give us the clearest method to discuss and explain the underlying vulnerability (secret-dependent control flow).
  2. 2.
    More complex networks or types of networks do not inherently offer greater security unless the underlying problem is addressed, even if demonstrating the attack might be more technically challenging.
  3. 3.
    The attack is agnostic to the accuracy of the network.
Victim Enclave
The enclave code runs inside of SGX. Given the constraints of SGX’s standard library replacement tlibc, we decided to employ Tensorflow Microlite, a software package for running (converted) Tensorflow lite models on microcontrollers, instead of “full” Tensorflow (which relies on a variety of system calls not easily provided by SGX). Our SGX enclave thus included a statically compiled Tensorflow Microlite library to perform inference. We converted our tensorflow models to tflite and then further exported them as to a byte array which can be compiled directly into a C program. These byte arrays are added to the enclave code which allows us to perform “secure” inference inside SGX. We note that for practical deployments of SGX-protected ML, using such a reduced variant of Tensorflow would ease deployment, reducing code and thus Trusted Computing Base (TCB) size, while maintaining compatibility with Tensorflow-trained models. The underlying implementation [sun_microsystems_tlibcs_1993] of exp() used by our network during inference comes from SGX’s libc alternative, tlibc.
Attack Code
To practically execute our attack, we used SGX-Step to collect the CPU instruction counts necessary to understand the state of our victim networks and apply the steps described above. Using SGX-Step, we developed an attack application, which starts the enclave and interacts with it. We use this to record page accesses which we later interpret into activation function states. Using SGX-Step, we log all page accesses within the enclave (after a certain trigger page was accessed) and interpret the results as part of a post-processing step.
Post-Processing
Once Tensorflow Microlite ran within an enclave, we profiled the execution and observed call traces (via sgx-gdb) to understand execution flow. An example can be seen in LABEL:lst:sgx_expf_stack. This lets us build an intuition for how Tensorflow Microlite performed inference and which functions and pages are accessed.
10x7ffff762b460 in expf()
20x7ffff7625e9c in tflite::(anonymous namespace)::ExpEval()
30x7ffff762318b in tflite::MicroInterpreterGraph::InvokeSubgraph()
40x7ffff760f04e in performInference()
50x7ffff7610577 in entry_point()
60x7ffff7611633 in sgx_entry_point()
70x7ffff76130a2 in do_ecall()
80x7ffff762c5a5 in enter_enclave()
90x7ffff762c7d5 in enclave_entry()
100x7ffff7f95309 in __morestack()
110x7ffff7f986fb in CEnclave::ecall()
120x7ffff7f9a852 in _sgx_ecall()
130x55555555bac9 in entry_point()
140x555555557a24 in main()
Listing 2: Simplified call stack for expf()-based neuron in Tensorflow Microlite.
Unlike sgx-gdb, SGX-Step only offers a relatively coarse-grained spatial resolution of 4 KiB memory pages, which prevents production enclave attackers from directly identifying the specific symbol (e.g., function or data) accessed by the enclave code. However, SGX-Step’s precise instruction-level temporal resolution enables the annotation of coarse-grained page-access traces with the exact number of instructions executed on each page. As illustrated in Figure 4, this capability allows to accurately reconstruct the inference execution structure, even without knowledge of the lower 12-bit page offset of the secret enclave instruction pointer (ERIP). By comparing the dotted red line which depicts the true ERIP values available to sgx-gdb and the black line which shows the reduced page-level granularity, one can see how little is lost. It is worth nothing that each page can contain multiple functions and so it is difficult to assign a precise “meaning” to each page without context. The instructions executed on one page could pertain to two very different activities at different points in execution. However, the combination of coarse-grained page-access patterns along with the exact amount of instructions executed per page, annotated in black text, clearly suffices to identify individual function calls or data accesses. In practice, this means that even if multiple symbols are co-located on the same 4 KiB page, we can still interpret the logs. We also know precisely the length (or length range) of different runs at different points during the execution, i.e., we can also tell if a trace is invalid and immediately reject it.
Figure 4: Visualization of SGX-Step’s instruction-granular page access trace. Inset of one layer’s 16 neurons. Note that the intra-page instruction pointer (RIP) values highlighted in dotted red are only displayed for reference visualization purposes here and are not used by the final attack.
By studying inference traces from the perspective of both sgx-gdb and SGX-Step (raw traces or visualizations like those shown in Figure 4), clear patterns emerge. For example, the function ExpEval() is located on a different page to expf() (each 0x1000 bytes is a new page) and we can plainly interpret the movement between them as the higher level neuron code invoking the lower level expf() math function 16 times (once for each neuron in this layer). The 18 instruction count long runs at the extreme left and right on page 0x2c000 seem to correspond to the start and end of a layer, which makes parsing each layer trivial. We note that the step counts provided by sgx-gdb and SGX-Step may vary slightly (but deterministically) due to the documented effects of macro-op fusion [moghimi2020copycat].
The final part of our tooling is a Python script and library which further abstracts and manages invocations of the enclave program. Because each neuron is accessed in order and each return occurs in sequence, it should be possible to monitor each neuron’s step count independently. In order to orchestrate the full attack, we use this script to interpret the log output by SGX-Step into a per-neuron state. Even though we receive the state for every neuron in each log, we only focus on one at a time. An example output of the running attack is shown in LABEL:lst:recovery_attack_output in Appendix A.
6.2 Multi-Layer Model with sigmoid()
Model
We now focus on a “multiplication” network (trained to multiply numbers), a small network with the following architecture: there are two inputs, followed by the first hidden layer which contains 4 sigmoid neurons. Then there is a second hidden layer containing 8 sigmoid neurons followed finally by the final layer with a single relu neuron.
Recovery
We follow the steps described in Section 5.2 and Section 6.1 to recover the first layer using the exact same attack code. The only difference is that we exploit a different CPU instruction count leakage; instead of targeting tlibc’s exponential function, we now attack the Tensorflow lite framework-level sigmoid function as described in Section 4.1. This shows how our attack is adaptable to different and more realistic activation functions, as well as side channel leakages elsewhere in the ML inference pipeline, cf. Section 7.1.1.
7 Results
We integrated our victim enclave, performing inference using our Tensorflow Microlite network, with SGX-Step for side-channel observations and our attack script for weight and bias recovery.
In Section 7.1 we discuss the results from the attack against the proof of concept regression network. In Section 7.2 we discuss how our findings relate to other activation functions in other ML Frameworks. As a part of that, we more explicitly consider relu() (Section 7.2.1) and framework code itself (Section 7.2.2).
7.1 End-to-End Attack Case Study for Tensorflow Microlite Enclave
Our attack was able to fully recover all the weights and biases for the neurons in the first layer of a Tensorflow Microlite network running inside an SGX enclave. After calibration, the full attack took 55 binary searches of the network per parameter to recover each of the 1200 parameter in the first layer down to or beyond 3 decimal places of accuracy.
During the attack, each first-layer neurons’ parameters were recovered at around 99% accuracy (when compared to the ground truth values) by <36% of the way through the full search (around search depth 20). The average error is <1% at this point. The error continues to slowly improve until the search hits the limits of a 32-bit floating point number at around depth 55. These findings are further summarised in Table II, which shows how the achieved average and maximum error in the recovered weights and biases varies with the number of queries per neuron in our 11 input regression ‘victim’.
The maximum error rates were higher in our full attack on the regression network than we had seen in smaller scale testing. We speculate this is due to numerical precision issues when solving certain neurons, particularly ones with very small weights or perhaps a combination of very low and very high magnitude weights. During the binary search there may be an ideal magnitude for the non-dynamic input values based on the magnitude of the dynamic input. We consider this an area for future study and discuss a workaround below.
The speed (e.g., number of invocations) is dependent on the architecture of the network (the number of parameters to recover) but also some factors the attacker can control such as the depth of the binary search. For example, the attack can be sped-up by only binary searching to a certain depth of iterations. As can be seen in Table II, we already achieve an < 1% average error rate by depth 20. Based on the desired level of accuracy, an attacker can choose to only search to some depth, rather than letting the binary search finish. We note that achieving sub 1% precision may not be as necessary as it seems. As the widespread adoption of quantization has shown, some models are still accurate with the precision loss due to the conversion to 8 bit integers.
In general, the total number $Q$ of queries to recover the first layer can be computed as:
$Q=\,P\,\cdot\,N\,\cdot S$
where $P$ is the number of hidden parameters (in the first layer, e.g., one weight per input plus one for the bias) to solve for, $N$ is the number of neurons in the first layer, and $S$ is the maximum number of steps in the binary search.
The $P$ value correlates to the minimum number of equations we require to solve a system of equations: one for each unknown. We can compensate for noise or other errors by adding extra equations to produce better solutions at the cost of more queries. In testing, finding 3 additional convergence sets helped reduce the maximum error by 28% at depth 25 compared to an attack where precisely $P$ convergence sets were generated (this is visualised in Figure 5 in Appendix C). It might be possible with future study to detect particularly weak systems of equations analytically and automatically add extra convergence sets to improve the results dynamically.
Secondly, we note the effect Calibration (Section 5.2.1) has on the performance of the attack. We can ensure we start the binary search in an overflow/underflow state by choosing a very high magnitude starting value for our dynamic input at the cost of having to waste many queries binary searching down to the range most hidden parameters occupy near 0. The performance of $S$ is therefore optimised by starting the search from the smallest possible point that still overflows our target. Rather than starting from an extreme value like MAX_FLOAT, a better alternative is to apply the heuristic that a weight likely has a magnitude between -1 and 1 and see if 100 is enough to overflow it. If not, we try 1000, then 10,000 and so on. By automating this calibration process for each neuron, we ensure a search with the fewest steps possible.
In our experiments we implemented calibration as a per-neuron process but we realised that this step could be performed across an entire layer simultaneously. In contrast to search phases of the attack, we can leverage the fact that we receive an instruction count measurement for all neurons simultaneously for each execution of the network. We create an input to the network that is all 0s and then iterate through each index in the input, replacing that 0 with increasing powers of 10 until we’ve under/overflowed the smallest weighted neuron in the layer. This maximal power of 10, $D$, when multiplied by $P-1$ (the number of inputs) allows us to calculate $C$, the number of executions required to calibrate the whole network.
$C=\,(P-1)\,\cdot D$
We note that $C$ is totally orthogonal to $N$ and $S$. Because $Q$ would likely be dominated by $N$ in a real network, this calibration phase should account for only a small fraction of the total queries in the full attack. By adding $C$ to $Q$ we get a good approximation for the number of executions required to recover the first layer.
TABLE II: Average and Maximum Error for different search iterations of 11 input Regression Network.
Iterations Average error Max. error
5 6.029863100281893 844.0875066255469
10 2.8364600666971387 786.2672913954896
15 0.27165280197604547 156.03829543970028
20 0.008040720460419709 4.501496529005743
25 0.0005365116007451519 0.25501229735941244
30 0.0003891028780509273 0.1506828257822338
35 0.00038239633382359514 0.14650184342297062
40 0.000382652272853835 0.14664602195446264
45 0.00038265156035156687 0.14664546962120745
50 0.000382651519816129 0.14664550955005118
55 0.00042219596795115005 0.14664550840498652
[…]
7.1.1 Results for Multiplication network
This recovery is summarised in Table III and is comparable to the results of our attack against the other model.
We are also able to recover partial (signless) convergence sets for the neurons in the second layer. For example, we can recover 35 sets for the first neuron in the second layer. Given its 4 inputs and single bias term (5 total unknowns) these are many more convergence point sets than we would need for a solution. However, because we do not have sign information, we are unable to conclusively solve for the neurons’ parameters.
TABLE III: Average and Maximum Error for different search iterations of 2-input Multiplication Network.
Iterations Average error Max. error
5 5.104921006404077 17.916865999472293
10 1.6756457515509167 17.51883507815715
15 1.5295341684026322 17.520958313634686
20 0.06962602190494313 0.7988771642785077
25 0.002053382484092837 0.023557097950552608
30 4.133456316032444e-05 0.0004519024401035132
35 3.313471981702896e-05 0.00036192843021964904
40 3.2159960239434036e-05 0.0003506233064154429
45 3.215762491123297e-05 0.00035059400594272816
50 3.215754896549253e-05 0.0003505930825000725
55 3.21575691725288e-05 0.0003505933114027471
7.2 Activation Function Safety Ecosystem Study
In order to extend our findings and consider the breadth of the issue, we examined implementations of common activation functions to determine in popular ML frameworks to assess if they exhibit input-dependent memory access patterns. For this, we used a combination of static (Ghidra) and dynamic (gdb) analysis to explore the often complex call stacks of these frameworks until an underlying math function was reached. Our findings are summarised in Table IV.
There are several dimensions to our survey: first, we found that in all cases we tested, the low-level math functions that underpin activation functions are not included in the ML frameworks themselves. Instead, they rely either on additional high performance math libraries, e.g., sleef in the case of PyTorch, Eigen and DNNL in the case of Tensorflow and XNN in the case of Tensorflow Lite or call into standard libraries like tlibc (in the case of SGX) or glibc on a Linux system. In some cases, these frameworks do both for different functions, adding to the complexity of the analysis.
The (in)security of an activation function can stem from the underlying library and implementation of these functions. For example, a call to sigmoid() that ends up in glibc’s highly input-dependent exp() in libm.so exposes the caller’s ML framework to vulnerability. Similarly, a call to the same function in a different framework that ends up in a highly optimised math library like Eigen may be harder to exploit. However, as we note in Section 4.1 and Section 6.2, higher level framework code can also be the culprit of exploitable memory access patterns. The hardware context may also play a roll in vulnerability; an Internet of Things (IoT) device performing inference on an embedded system for example may not have the hardware functionality or enough space for advanced math routines beyond a basic standard library.
The reason optimised math libraries may be harder to attack is that their use of vectorised instructions (for example AVX on Intel or NEON on Advanced RISC Machines (ARM)) which discourage branching (since that can affect performance). For example, we observed that in PyTorch, softmax() was realised through calls into the sleef library, which at default optimisations compiled into AVX assembly code without memory access leakage. We note however that sleef incorporates a dispatch mechanism that calls into different implementations depending on the outcome of cpuid. Notably, if this library was used in an SGX enclave, where cpuid is typically implemented as an untrusted ocall [^3] https://github.com/intel/linux-sgx/blob/80a662/common/inc/sgx_tstdc.edl, the adversary could spoof the cpuid result and force usage of the insecure non-AVX version. Curiously, for sigmoid(), PyTorch resorts to the standard library implementation of exp(), rendering the implementation insecure independent of cpuid. Tensorflow uses Eigen for all the functions we tested except for Exponential(), which calls directly into glibc. Finally it is worth restating that even though vectorised code is non-branching, it still may not be constant time, which we discuss further in Section 8.3.
Finally, Table IV shows how the optimisation level has a direct impact on the security of relu() in the case of Tensorflow Microlite as further discussed in Section 7.2.1. While the compiler default of -Os happens to be free of memory access leakage, this is more by accident than by design, which is not desirable in the case of libraries that handle sensitive data. Analogously, the security benefits (with regard to the attack we have demonstrated) of vectorised code is not a proactive security-related decision but a side effect of pursuing performance.
TABLE IV: Survey of vulnerability across various ML and standard libraries and for different optimisation levels.
ML Framework RELU Exponential Sigmoid Softmax
TFLiteMicro (SGX/tlibc, -O0)
TFLiteMicro (SGX/tlibc, -Os)
TFLiteMicro (glibc, -Os)
TFLite (glibc)
TensorFlow CPU (glibc)
PyTorch (glibc) n.a. ✓*
Legend: ✓: Secure ✗: Insecure due to data-dependent access pattern
*: Only secure if cpuid indicates AVX support
7.2.1 Vulnerability of relu()
In addition to expf(), we also considered relu() and its implementation in Tensorflow Microlite, which relies on the std::max() function. When running inside of an enclave, the std::max() library call is provided by an implementation in SGXs tlibc library which can be seen in LABEL:lst:max_in_c).
1template <class _T1>
2struct __less<_T1, _T1>
3{
4 _LIBCPP_INLINE_VISIBILITY _LIBCPP_CONSTEXPR_AFTER_CXX11
5 bool operator()(const _T1& __x, const _T1& __y) const {return __x < __y;}
6};
7
8template <class _Tp, class _Compare>
9inline _LIBCPP_INLINE_VISIBILITY _LIBCPP_CONSTEXPR_AFTER_CXX11
10const _Tp&
11max(const _Tp& __a, const _Tp& __b, _Compare __comp)
12{
13 return __comp(__a, __b) ? __b : __a;
14}
Listing 3: std::max() source code from sgxsdk/include/libcxx/algorithm.
With the default settings (-Os passed to the compiler, gcc 11.4.0, and the function marked as inline), the above code was compiled into branchless (e.g., UCOMISS) instructions as shown in LABEL:lst:max_in_asm. In contrast, when compiling with -O0, the function was assembled with conditional branches as shown in LABEL:lst:max_in_asm_o0. This is significant because when the compiler emitted jmp-class instructions, it was possible (using sgx-gdb) to detect the single step difference and therefore discern between the cases where $input\leq 0$ and $input>0$. This is a threshold point between two instruction count classes about $0.0$, similar to the ones discussed previously in expf(). Though the difference might only be a single CPU instruction (whether a jump was taken or not) this difference is detectable.
100100a04 MOVSS XMM0,dword ptr [param_2]
200100a08 UCOMISS XMM0,dword ptr [param_1]
300100a0b MOV RAX,param_1
400100a0e CMOVA RAX,param_2
500100a12 RET
Listing 4: Resulting assembly for std::max() at -Os with gcc 11.4.0.
1001019c9 PUSH RBP
2001019ca MOV RBP,RSP
3001019cd MOV qword ptr [RBP + local_10],param_1
4001019d1 MOV qword ptr [RBP + local_18],param_2
5001019d5 MOV RAX,qword ptr [RBP + local_10]
6001019d9 MOVSS XMM1,dword ptr [RAX]
7001019dd MOV RAX,qword ptr [RBP + local_18]
8001019e1 MOVSS XMM0,dword ptr [RAX]
9001019e5 COMISS XMM0,XMM1
10001019e8 JBE 001019f0
11001019ea MOV RAX,qword ptr [RBP + local_18]
12001019ee JMP 001019f4
13001019f0 MOV RAX,qword ptr [RBP + local_10]
14001019f4 POP RBP
15001019f5 RET
Listing 5: Resulting assembly for std::max() at -O0 with gcc 11.4.0.
Within the our case study, relu() within Tensorflow Microlite was not vulnerable, but it was not secure by design either. This is worrisome because in contrast to vetted cryptographic code, ML libraries currently do not exhibit input-independent memory access and execution behavior, and further the eventual security properties of high-level ML code may depend on compiler optimizations and other non-explicit behavior as with relu() in some configurations. Without any specific safeguards taken, ML libraries are not intrinsically secured against the vulnerability discussed in this paper.
As a sidenote, though not considered in this paper, we note that on x86, many floating point instructions can raise exceptions or otherwise exhibit operand-dependent timing as mentioned by [Gongye20] and (in a different context) [Kohlbrenner17], which might result in timing vulnerabilities even if branchless code is emitted. We leave the exact measurement and exploitation of such timing side channels in real-world settings as an area for future research.
7.2.2 Vulnerability of Framework Code
As discussed in Section 4.1, and demonstrated practically in Section 5.3, exploitable memory access pattern vulnerabilities can exist outside mathematical library code in ML frameworks themselves. This further supports our case that these frameworks are not suitable for enclave execution under a Side-Channel Analysis (SCA) advisory. Beyond the security properties of the libraries they use, security-conscious ML framework developers need to also consider their own code and if they are introducing exploitable timing side channels. This is especially true for more complex activation functions like softmax() which are not common mathematical functions likely to be found in standard math libraries.
8 Discussion and Mitigations
First, we demonstrated that SGX-protected ML implementations can exhibit vulnerabilities due to oversights in their implementation, leading to extraction of precise model weights and biases. Our proof-of-concept shows that an attacker with a high-resolution controlled or side channel is able to exploit data-dependent memory accesses to recover model details, whereas prior software side channel attacks typically only recovered hyperparameters such as model architecture [yan_cache_2020]. Though there are certain limitations, precise weight recovery can be performed with at maximum 55 queries per weight, and at over 99% accuracy at less than half that number (plus calibration). Speed can be parametrically adjusted to account for the desired accuracy.
While our attack demonstration focused on the popular Intel SGX architecture, the underlying insights and weight-recovery techniques are applicable to other TEEs. Notably, the required side-channel primitives offered by SGX-Step, i.e., precise single-stepping and page accesses, have been demonstrated on alternative VM-based TEEs like AMD SEV [Wilke_Wichelmann_Rabich_Eisenbarth_2023] and Intel TDX [Wilke24, shome2024tdexit-notify].
Ultimately, while not in the scope of this paper, we believe that if an attacker can obtain precise-enough timing measurements, similar attacks may be possible outside of a TEE context, e.g., for virtualised cloud deployments. Cache attacks with high temporal resolution, e.g., Prime+Scope [Purnal21], could be applicable to recover the access patterns similar to SGX-Step, though attacks would likely require modifications to deal with increased noise.
We further show that input-dependent memory accesses and branching are prevalent in ML implementations, and that the necessity for secure programming practices is not widely understood in the ML community.
Even if the attack described in this paper currently does not apply to non-enclaved targets, we argue that ML library developers should pay attention to the implementation attack surface. For example, even if relu() implementations were often secure in the surveyed libraries, this largely relied on certain compiler optimizations and was not ‘by design’. Similar to cryptographic libraries, ML libraries should hence consider hardening their codebase against software side channels such as data-dependent memory access patterns.
Security and privacy-sensitive algorithms, to which ML now belongs, must be robust by design, regardless of context, platform, or (mis)use of library interfaces by a programmer.
8.1 Countermeasures
As with the work done in cryptography libraries, there are established steps to avoid data-dependent branching and memory accesses, for example by relying on bitwise logical operations, constant-time conditional moves, and other related techniques. For x86, as discussed for the sleef library, using vectorised SIMD code can also mitigate the issue, though needs to be thoroughly vetted for other side channels like floating point timing [Kohlbrenner17, Gongye20].
At present, akin to a “chosen plaintext” attack for cryptographic algorithms, our attacks entails sending arbitrarily large or small floating point numbers into the network. Any input normalization or quantisation might mitigate the fast calibration and binary search approaches described in this paper, only enabling partial recovery through the grid search technique. Similarly, as the attack requires a large number of queries to the network, rate-limiting or detection of “suspicious” usage would render exploitation harder in practice. However, again, ML security should in our view not rest on developers having to normalise inputs or limit performance, so these measures are supplementary to a securely developed library.
Finally, researchers have explored defenses at the level of the TEE itself to not fully eliminate side-channel leakage, but frustrate exploitation in practice. Initial approaches [shih2017tsgx, gridin2020dejavu, oleksenko2018varys] focused on detecting suspicious interrupt rates as a side-effect of an ongoing controlled-channel attack, which may be error-prone and suffer from false positives. Alternatively, custom oblivious RAM solutions [aga2019invisipage, zhang2020klotski, brasser2019dr] may probabilistically hide secret-dependent enclave memory accesses, however this is at the cost of prohibitive performance overheads. Most prominently, AEX-Notify [constable2023aexnotify] is a recent Intel SGX hardware-software extension that aims to eliminate SGX-Step’s single-stepping capabilities by prefetching selected application pages. Notably, AEX-Notify is an opt-in feature that is only available in recent CPUs. With regards to our attacks, victim enclaves that opt-in to AEX-Notify would preclude the ability to gather precise instruction counts through single-stepping. However, AEX-Notify explicitly does not prevent general information leakage through page faults and, hence, cannot fully mitigate our attacks.
8.2 Limitations
There are a certain limitations to our proof of concept.
We require the victim to run in a co-located TEE in order to use our controlled channel, which is a different requirement than previous model stealing attacks, which require chosen inputs and the output distribution of the inference pass.
Despite the recovery of the first layer of neurons fundamentally undermining SGX’s promise to secure the entire computation, our attack is limited in recovering parameters from deeper layers, depending on the model architecture. This is mostly due to the (limited) expressive power of some common activation functions as noted in Table I and detailed in Section 5.3. Besides, the used grid search is less efficient than binary search for the first layers, and depending on the magnitude of the weights, small search steps might be needed to discover thresholds. An adaptive approach could be used where the sweep region is deepened until a threshold is hit. We note however that small weights likely contribute little to the network, and thus could be replaced by a static value during the recovery process, producing an equivalent (though not identical) network.
8.3 Future Work
To efficiently recover missing parameters from deeper layers, if the output of the network is available, our attack could be combined with a Tramèr-style attack. After a partial parameter recovery, the obtained parameters would be placed into an untrained network with identical architecture, and be locked during training as in transfer learning. Then, the remaining neurons can be trained using examples generated from input/output pairs from the source network. In particular, image processing and similar networks contain a large proportion of their parameters in the early layers (and have an overall funnel shape), hence, this might result in a more precise recovery compared to a brute-force approach.
Besides, we demonstrated our attack approach for simple FNNs. It would be interesting to expanding the attack to more complex architectures like Convolutional Neural Networks or transformers, using a library OS like Gramine [^4] https://github.com/gramineproject/gramine to run the relevant code in an SGX enclave. To improve the performance of our attack, a weighted binary search could reduce the number of search steps, as could caching some shared state such as the during the calibration step.
Finally, we focused on input-dependent memory access patterns and instruction counts in this paper, but did not take input-dependent latency of single instructions into account.
For example, the UCOMISS instruction discussed in Section 7.2.1 is not necessarily guaranteed to execute in constant time [Gongye20], and can raise floating point exceptions. This could render vectorised code like the implementation from sleef vulnerable to a modified version of our attack: while different inputs might have identical instruction counts (as reported by SGX-Step), the actual runtime might be different and be observable with other side channels like IRQ latency [VanBulck18].
9 Conclusion
The recent surge in ML applications necessitates an in-depth understanding of new types of attack vectors and defenses. In this respect, confidential computing architectures, and in particular Intel SGX, have gained significant traction in recent years and hold the potential to securely outsourcing critical ML computations to pervasive untrusted remote cloud platforms. This paper showed, however, that SGX-protected ML implementations exhibit subtle side-channel vulnerabilities that allow to extract precise model weight and biases in practical proof-of-concept attacks. In the wider perspective, our work may generalise to other TEEs and highlights the issue of input-dependent memory accesses that are prevalent in today’s ML implementations, suggesting that security-critical ML applications should consider adopting coding practices studied in the cryptographic community.
Acknowledgments
The PhD of Jesse Spielman is partially funded by a donation from Intel. This research was further partially funded by the Engineering and Physical Sciences Research Council (EPSRC) under grants EP/X03738X/1 and EP/R012598/1. This research is partially funded by the Research Fund KU Leuven, and by the Cybersecurity Research Program Flanders.
[…]
Appendix A Example Attack Output
1 [ Neuron 2 ]
2 stats:
3 952.99 seconds since start
4 1439 executions so far
5 4 Incomplete logs
6 0 non-deterministic steps
7 6 neuron_check failures
8 Function = exp
9 Search strategy: seeded binary search
10 Calibrating:
11 count: 14
12 recovered signs: + - - - + - - - - + +
13 recovered maxvals: 1000 1000 1000 1000 1000 10000 1000 1000 10000 10000 1000
14 Finding convergence points for equation 1
15 depth 0: Overflow
16 depth 10: Normal
17 depth 20: Overflow
18 depth 30: Normal
19 depth 40: Overflow
20 depth 50: Overflow
21 Finding convergence points for equation 2
22 <SNIP>
23 Finding convergence points for equation 12
24 depth 0: Overflow
25 depth 10: Normal
26 depth 20: Normal
27 depth 30: Normal
28 depth 40: Overflow
29 depth 50: Overflow
30 deepest solution: [
31 0.14825567 -0.17573831 -0.28457861 -0.24754734
32 0.09355622 -0.09359772 -0.2866397 -0.2083804
33 -0.10068617 0.07542896 0.10219476 -0.04099777]
34 ground truth: [
35 0.14825566 -0.17573832 -0.28457862 -0.24754737
36 0.09355621 -0.09359772 -0.28663969 -0.2083804
37 -0.10068618 0.07542896 0.10219476 -0.04099637]
38 abs percent error: [
39 0.0%, 0.0%, 0.0%, 0.0%, 0.0%, 0.0%,
40 0.0%, 0.0%, 0.0%, 0.0%, 0.0%, 0.0%]
41 Checkpoint saved!
Listing 6: Example output from our attack tool finding convergence points.
Appendix B Implementation of sigmoid() (called logistic()) in Tensorflow lite
1inline void Logistic(const RuntimeShape& input_shape, const float* input_data,
2 const RuntimeShape& output_shape, float* output_data) {
3 const float cutoff_upper = 16.619047164916992188f;
4 const float cutoff_lower = -9.f;
5
6 const int flat_size = MatchingFlatSize(input_shape, output_shape);
7
8 // [comments removed]
9
10 for (int i = 0; i < flat_size; i++) {
11 float val = input_data[i];
12 float result;
13 if (val > cutoff_upper) {
14 result = 1.0f;
15 } else if (val < cutoff_lower) {
16 result = std::exp(val);
17 } else {
18 result = 1.f / (1.f + std::exp(-val));
19 }
20 output_data[i] = result;
21 }
22}
Listing 7: Implementation of Logistic() in Tensorflow lite
Appendix C Error Comparison with Extra Convergence Sets
Figure 5: Note the effect that extra convergence sets have to reduce variance and lessen the errors, especially after depth=25

### v2

@@
-bibliography.bib
-[…]
-bibliography.bib
+Triple DES AES Advanced Encryption Standard
 3DES
@@
 Non-Linear Function
+NN
+Neural Network
 NFC
@@
 Operating System
-VM
-Virtual Machine
 ML
@@
 Memory-Mapped I/O
-Activation Functions Considered Harmful: Recovering Neural Network Weights through Controlled Channels
-Jesse Spielman and David Oswald and Mark Ryan and Jo Van Bulck School of Computer Science
-University of Birmingham
-Birmingham, UK
-jxs1366@student.bham.ac.uk
-School of Computer Science
-University of Birmingham
-Birmingham, UK
-d.f.oswald@bham.ac.uk
-School of Computer Science
-University of Birmingham
-Birmingham, UK
-m.d.ryan@bham.ac.uk
-DistriNet
-KU Leuven, Belgium
-Leuven, Belgium
-jo.vanbulck@cs.kuleuven.be
+Activation Functions Considered Harmful:
+Recovering Neural Network Weights through Controlled Channels
+Jesse Spielman and David Oswald and Mark Ryan and Jo Van Bulck
 Abstract
-With high-stakes machine learning applications increasingly moving to untrusted end-user or cloud environments, safeguarding pre-trained model parameters becomes essential for protecting intellectual property and user privacy.
-Recent advancements in hardware-isolated enclaves, notably Intel SGX, hold the promise to secure the internal state of machine learning applications even against compromised operating systems.
-However, we show that privileged software adversaries can exploit input-dependent memory access patterns in common neural network activation functions to extract secret weights and biases from an SGX enclave.
-Our attack leverages the SGX-Step framework to obtain a noise-free, instruction-granular page-access trace.
-In a case study of an 11-input regression network using the Tensorflow Microlite library, we demonstrate complete recovery of all first-layer weights and biases, as well as partial recovery of parameters from deeper layers under specific conditions.
-Our novel attack technique requires only 20 queries per input per weight to obtain all first-layer weights and biases with an average absolute error of less than 1%, improving over prior model stealing attacks.
-Additionally, a broader ecosystem analysis reveals the widespread use of activation functions with input-dependent memory access patterns in popular machine learning frameworks (either directly or via underlying math libraries).
+Recent advancements in hardware-based enclaved execution environments, such as Intel SGX, aim to protect critical model parameters in high-stakes machine learning applications increasingly moving to end-user or cloud environments.
+However, this also introduces the risk of privileged side-channel attacks, traditionally aimed mainly at cryptographic targets.
+In this paper, we develop a novel attack methodology that exploits input-dependent memory access patterns in common neural network activation functions to extract hidden model parameters.
+In several case studies using the SGX-Step attack framework and the Tensorflow Microlite library, we demonstrate complete recovery of first-layer weights and biases, along with partial recovery of deeper layer parameters under specific conditions.
+Our novel attack technique requires only 20 queries per input per weight to obtain all first-layer weights and biases, with an average absolute error of less than 1%, improving over prior model stealing attacks.
+Furthermore, a broader ecosystem analysis reveals the widespread use of activation functions with input-dependent memory access patterns in popular machine learning frameworks and maths libraries.
 Our findings highlight the limitations of deploying confidential models in SGX enclaves and emphasise the need for stricter side-channel validation of machine learning implementations, akin to the vetting efforts applied to secure cryptographic libraries.
-Index Terms:
-Machine Learning Security, Side Channel Attacks, Trusted Execution, Confidential Computing, SGX
-1 Introduction
+I Introduction
 Training and deploying Machine Learning (ML) models, in particular Neural Networks, often requires a massive investment in computational resources.
@@
 An application that incurs a fee to access per interaction could also be duplicated to remove this restriction.
-While existing model stealing attacks show how to duplicate neural networks essentially via brute force, they scale poorly for increasingly complex networks and require detailed access to output probabilities or physical access for measurements such as power consumption.
+While existing model stealing attacks show how to duplicate NNs essentially via brute force, they scale poorly for increasingly complex networks and require detailed access to output probabilities or physical access for measurements such as power consumption.
 Because ML (both training and inference) requires substantial computation, there is a strong focus on performance.
-Further, because of the hardware requirements of ML training and inference (such as in the case of an AI model which backs an end user-accessible web interface), there is an increasing need to host ML workloads in the cloud, e.g., for scaling or to use the latest Central Processing Unit (CPU) and Graphics Processing Unit (GPU) hardware.
-However, this move off-premises and onto (often shared) hardware introduces a number of security concerns.
-To mitigate the risk, substantial work has been done using Trusted Execution Environments such as Intel Software Guard Extensions (SGX) or AMD Secure Encrypted Virtualization (SEV) to protect ML workloads running on third-party hardware [russinovich2023confidential, intel_secure_ml, kumar23, li2024coreguard, kethireddy2024secure].
-A core component of neural networks are activation functions, chosen by developers (alongside other architectural decisions such as the number of layers or the learning rate) to introduce non-linearity.
+Further, because of the hardware requirements of ML training and inference (such as in the case of an AI model which backs an end user-accessible web interface), there is an increasing need to host ML workloads in the cloud.
+However, this move onto (often shared) hardware introduces a number of security concerns.
+To mitigate the risk, substantial work has been done using hardware-based Trusted Execution Environments such as Intel Software Guard Extensions (SGX) or AMD Secure Encrypted Virtualization (SEV) to protect confidential ML workloads [1, 2, 3, 4, 5].
+These TEEs are presumed to facilitate the secure and efficient outsourcing of critical ML computations to third-party hardware, while simultaneously preserving the confidentiality of model parameters that embody key Intellectual Property (IP).
+A core component of NNs are their activation functions, chosen by developers (alongside other architectural decisions such as the number of layers or the learning rate) to introduce non-linearity.
 These functions are generally computed once for each neuron based on the sum of the multiplication of all incoming inputs by their weights (plus a bias term).
-Often, these functions are provided by a ML framework such as Tensorflow or PyTorch, which all include routines for training, running inference, and deploying models.
-It is important to note that these framework-provided activation functions usually rely on basic mathematical functions such as std::max() or exp(), often provided by the system’s standard libraries (e.g., glibc or musl) and optimised for speed, not side-channel resistance.
+Often, these functions are provided by a ML framework such as Tensorflow or PyTorch, which include routines for training, running inference, and deploying models.
+It is important to note that these framework-provided activation functions usually rely on basic mathematical functions such as std::max() or exp(), which are often provided by the system’s standard libraries (e.g., glibc) and optimised for speed, not side-channel resistance.
 TEEs like Intel SGX often supply their own (restricted) standard library (e.g., SGX’s tlibc), which provides implementations known to work within the special enclaved environment.
-In summary, ML frameworks inherit their security properties from the used standard library math functions.
-Research into side-channel attacks (both with physical access as well as remotely exploitable side channels like timing) has led to the “hardening” of certain security sensitive algorithms, especially in cryptographic libraries.
+Thus, ML frameworks inherit some security properties from underlying standard library maths functions.
+Research into Side-Channel Analysis (SCA) (both with physical access as well as remotely exploitable side channels like timing) has led to the “hardening” of certain security sensitive algorithms, especially in cryptographic libraries.
 However, less attention has been paid to the side-channel security of ML libraries, even though their parameters (weights and biases) are today often akin to cryptographic secrets in sensitivity.
@@
 Particularly, we develop a novel methodology to recover partial weights and biases from the first (and deeper) layer(s) of a victim network when provided with an instruction-granular page-access trace.
-As a practical case-study attack, we deploy a Tensorflow Microlite model inside an SGX enclave and extract deterministic, instruction-granular page access traces using the SGX-Step [van_bulck_sgx-step_2017] framework.
-We demonstrate successful recovery of each weight and bias of the network to more than 5 decimal places accuracy with 55 invocations of the network, or with less than 1% average error with around 20 invocations after an initial calibration phase.
-Our end-to-end attack outperforms Tramèr-style model stealing attacks [tramer_stealing_2016], which require around 100 invocations for each parameter in simpler classification networks.
+As a practical case-study, we deploy three Tensorflow Microlite models inside SGX enclaves and extract deterministic, instruction-granular page access traces using the SGX-Step [6] framework.
+Our first case study demonstrates how we can use those traces to successful recovery all first layer weights and biases to more than 5 decimal places of accuracy with 55 invocations of the network, or with less than 1% average error with around 20 invocations after an initial calibration phase.
+Our end-to-end attack outperforms [7] model stealing attacks [7], which require around 100 invocations for each parameter in simpler classification networks.
+We then use our second and third case studies to explore deeper layers and much larger networks, respectfully.
 Considering the wider ecosystem, we survey widely used ML libraries and find widespread use of secret-dependent access patterns in activation functions.
-Furthermore, we discuss to what extent our attack vector is applicable to other popular TEE architectures.
-Contributions
-Summarised, our contributions are:
+Furthermore, we discuss to what extent our attack vector applies to other TEEs.
+Contributions. Summarised, our contributions are:
   • A novel methodology to recover weights and biases from partial memory-access side-channel traces.
-  • An end-to-end case study of practical memory-trace extraction and accurate weight/bias recovery from Tensorflow Microlite running inside an SGX enclave.
+  • Three end-to-end case studies demonstrating concretely how, by exploiting our memory-trace extraction technique, we can perform accurate weight/bias recovery from Tensorflow Microlite networks running inside SGX enclaves in a few different configurations.
   • A comprehensive survey of input-dependent accesses in common activation functions across popular ML and standard libraries.
   • An open-source Tensorflow Microlite SGX benchmark for future work on attacks and mitigations.
-Ethical Considerations
-This work contributes to better understanding limitations of widely used TEE technology.
-Our attacks exploit the well-documented issue of precise monitoring of enclave memory accesses using tools like SGX-Step.
-All experiments were conducted on proof-of-concept implementations on our own local machines.
-2 Background and Related Work
-Trusted Execution Environments
-One of most commonly used and widely studied TEEs is Intel’s SGX, which, even though discontinued for client CPUs, is widely supported on Xeon Scalable server CPUs and now marketed for use cases such as secure ML deployments [intel_secure_ml].
-SGX-secured code is executed in a hardware-protected ‘enclave’ (with its memory encrypted) such that even the root user cannot gain access.
-SGX does not make any guarantees about protections against timing and memory-based side channels [noauthor_intel_nodate], and hence is vulnerable to cache and page fault attacks [nilsson_survey_2020].
+Ethics and Open Science.
+All experiments were conducted on Proof-of-Concept (PoC) implementations on our own local machines.
+To ensure the reproducibility of our results, and to enable future research on side-channel attacks and defenses, we release all code and data as open-source at https://github.com/heavyimage/afch_paper.
+Scope.
+To the best of our knowledge, this is the first study to thoroughly analyse side-channel implications of deploying ML workloads in TEEs, which are increasingly being suggested by industry and academia to preserve IP and end-user privacy.
+In this paper, we do not claim a fully-fledged, optimised and weaponised real-world attack; instead we focus on developing a novel attack methodology to steal partial model parameters via case-study-driven evaluation.
+While full recovery of real-world networks may be out of reach, we show that motivated adversaries can fully automatically extract thousands of hidden, supposedly secure parameters.
+Following the rich tradition in cryptography, where a long line of increasingly potent side-channel attacks have highlighted the need for constant-time programming practices, we hope our work will draw attention to the security and privacy trade-offs of using TEEs for private ML and underline the importance of side-channel-resistant coding practices.
+II Background and Related Work
+Trusted Execution Environments.
+One of most widely studied TEEs is Intel’s SGX, which, even though discontinued for client CPUs, is widely supported on Xeon Scalable server Central Processing Units and now marketed for use cases such as secure ML deployments [2].
+SGX-secured code is executed in a hardware-protected enclave (with its memory encrypted) such that even the root user cannot gain access.
 SGX also allows secrets to be securely loaded into the enclave remotely, and subsequently be (un)sealed for local storage.
 This allows data of the enclave (such as a pre-trained ML model) to be also protected while at rest.
-Other TEEs of note include ARM’s TrustZone which brings the concept of TEE to e.g., mobile devices for isolating certain software from the user, for example, in Digital Rights Management (DRM) applications [arm_tz].
-Intel’s Trust Domain Extensions (TDX[intel_tdx] and AMD’s SEV [amd_sev] are TEEs that build on application-level TEEs which provide hardware-backed isolation guarantees at the level of entire Virtual Machines.
-Their approach is intended to make it easier for developers to take advantage of TEE security features, because applications do not need to be ported or modified for the enclaved security model.
-Though not a vulnerability per se, a consequence of SGX’s adversary model is that it allows an attacker to have precise control over the enclave’s execution.
-This gives rise to tools like SGX-Step [van_bulck_sgx-step_2017], which allows an attacker to perfectly single-step enclaved code by causing the enclave to be interrupted, suspended, and then resumed rapidly.
-When paired with a more conventional side channel that can probe memory pages to detect other accesses, a precise record of a program’s execution (or trace) with respect to those pages can be recorded.
-Model Stealing Attacks
-[…]
-Side-Channel Attacks on ML Implementations
-[…]
-[…]
-[…]
-[…]
-In contrast, side-channel attacks that recover the full model with physical access to the hardware and measurements of electro-magnetic (EM) emanation or power consumption have been further developed [Hua18, Batina19, Yu20, horvath2024barracuda].
+SGX does not provide guarantees against timing or memory-based side channels [8], rendering enclaves vulnerable to cache and page fault attacks [9, 10].
+Moreover, due to SGX’s privileged adversary model, such attacks can achieve high temporal resolution using tools like SGX-Step [6], which exploit timer interrupts to single-step enclave execution, effectively interleaving enclave execution with attacker-controlled code at an instruction-level granularity.
+Other TEEs of note include ARM’s TrustZone which brings the concept of TEE to mobile devices [11].
+Intel’s Trust Domain Extensions (TDX[12] and AMD’s SEV [13] are TEEs that provide hardware-backed isolation guarantees at the level of entire Virtual Machines.
+Model Stealing Attacks.
+The recovery of parameters from a trained neural network via “model stealing” is well documented in the literature.
+Such practical recovery attacks, first demonstrated by [7], enabled the duplication (stealing) of neural networks and other models [7].
+In these attacks, the original network is used as an oracle to train a duplicate network with similar hyperparameters.
+The attack works by choosing inputs via a search of the parameter space of the network (using one of several different heuristics) and joining that input with the corresponding output probabilities which are returned.
+These data pairs are used as training examples for the new network.
+“Solving” a network in this way results in the calculation of hidden parameters that are derived from but not necessarily precise to the original source network.
+Furthermore, they require a large number of queries, around 100 per parameter, as well as access to the output probabilities to essentially brute-force the space of the network.
+While these attacks may reconstruct the overall structure and general features of the network, it remains an open question how effective they are at reproducing the specifics of the original network, e.g., when attempting to transfer adversarial examples from duplicate to original.
+An alternative approach by [14] (and references therein) treat the recovery of parameters as cryptanalysis, showing that the sign information for each neuron in a 1.2M parameter CIFAR10 network can be recovered in around 30 min using a 40 GB A100 Graphics Processing Unit (GPU[14].
+Side-Channel Attacks on ML Implementations.
+An alternative approach to recover model parameters is to use a software-based (timing or cache) side channel.
+Most research in this direction has focused on the recovery of hyperparameters, such as the network’s architecture, but not the precise weights: [15] demonstrate a technique for recovering the hyperparameters of a victim network such as the number of layers and their activation functions using a Flush+Reload cache attack [15].
+Other works in this direction have used similar cache attacks [16], execution time [17], looked at embedded devices [18], and considered equivalent approaches on GPUs [19, 20, 21].
+Other uses for software-based SCA include the recovery of network inputs from floating point timing [22], membership inference [23], and the generation of adversarial examples [24, 25].
+Because the targeted leakages are much smaller and hence harder to exploit, less attention has been paid to full model recovery, i.e., weights and biases, through software-based side channels: [26] show how an attacker-controlled configuration of the x86 floating point unit can be used to recover weights from a toy NN (using a custom implementation) running in an SGX enclave [26].
+[27] show that, if an adversary obtains the precise timing of each network layer, in certain cases they can use timing leakages from floating point to recover the full model [27].
+In particular, the authors rely on different execution timing on certain x86 CPUs when processing “subnormal” numbers.
+[27] do not specify whether their attack applies to real-world ML libraries, and assume that the necessary precise timing measurements “can be achieved by either analysing the cache access pattern […] or through visual inspection of power traces”.
+In contrast to their work, in this paper, we consider the input-dependent memory access patterns of activation functions in widely used ML libraries, showing an end-to-end attack that uses SGX-Step for reliable side-channel observations.
+In contrast, side-channel attacks that recover the full model with physical access to the hardware and measurements of electro-magnetic (EM) emanation or power consumption have been further developed [28, 29, 30, 31].
 However, as these attacks require measurements on or in close proximity to the targeted chips, they are of less relevance to datacenter applications, where physical access to the underlying servers is typically closely guarded.
 Hence, in this paper, we focus on software-based side-channel attacks to recover weights/biases without physical access and without access to the output probabilities.
-Controlled Channel Attacks
-For SGX, Intel explicitly places the burden of preventing secret-dependent memory accesses for security-critical code on the developer [Wenhao17].
+Controlled Channel Attacks.
+For SGX, Intel explicitly places the burden of preventing secret-dependent memory accesses for security-critical code on the developer [32].
 In practice, this means that enclave developers should avoid data-dependent patterns of the form array[x] or if(x) func(), where x depends on a secret.
 It is important to note that in practice, such secret-dependent access patterns may only become evident at the machine code level, as optimising compilers might e.g., opt to compile “branchless” code into code with conditional branches, and vice versa.
-[…]
+Apart from well-known microarchitectural cache leakage [33], which is notoriously noisy, a number of SGX-specific side channels were discovered that exploit the privileged attacker’s control over the untrusted Operating System (OS).
+One of the first ones being “controlled channels” as introduced by [10] [10].
+This method exploits the fact that page faults within an SGX enclave are handled by the untrusted OS.
+A privileged adversary can, hence, temporarily unmap a page, observe whether a page fault occurs, and from this fully deterministically infer enclave memory access patterns at a 4 KiB, page-level spatial granularity.
+While such page-fault attacks have been proven particularly powerful, e.g., to extract text and images [10] or full cryptographic keys [34], their relatively coarse-grained spatial resolution may limit exploitability.
+Consider the example code snippet provided in Figure˜1, where a tight strlen loop is executed that fits entirely within a single code and data page.
+As subsequent accesses to the same page are cached in the processor’s Translation Lookaside Buffer (TLB), and the enclave needs both the code and data page to make forward progress, page-fault adversaries only observe the first access to a page and are not able to distinguish successive strlen loop iterations.
+To overcome this limitation, [35] showed that privileged adversaries may also monitor page-table attributes [35] using the ‘accessed’ and ‘dirty’ bits to count accesses to a certain page.
 Figure 1: Victim enclaves exhibiting tight memory-access patterns can be precisely interrupted at instruction-level granularity using SGX-Step allowing to extract deterministic page-access count traces.
-The open-source SGX-Step framework [van_bulck_sgx-step_2017, vanbulck2023sgxstep] allows for precise single-stepping of production enclaves using privileged x86 APIC timer interrupts, such that page access patterns can be deterministically monitored for every enclave instruction, giving precise insights into the operation of a victim enclave.
-By conservatively under-estimating the APIC timer interval, SGX-Step results in either zero or single-steps, but always avoids multi-steps [vanbulck2023sgxstep, Table 5].
-Subsequent works [VanBulck18, moghimi2020copycat, aldaya2020one, van_bulck_tale_2019, constable2023aexnotify] showed that the accessed (A) bit in the enclave code page deterministically distinguishes single-steps (enclave instruction retired; A=1) from zero-steps (A=0).
-Intel has acknowledged perfect single stepping in a recent paper [constable2023aexnotify], which also explains the underlying mechanism to stall the CPU pipeline using microcode assists.
-To date, instruction-granular page-access traces extracted with SGX-Step have been repeatedly abused to reliably exploit enclave interface vulnerabilities [van_bulck_tale_2019] or to reconstruct cryptographic key material [moghimi2020copycat, aldaya2020one].
-In the following, we show that such traces are sufficient to recover weights from real-world neural network libraries running inside a production SGX enclave.
-3 System and Adversary Model
+The open-source SGX-Step framework [6, 36] allows for precise single-stepping of production enclaves using privileged x86 APIC timer interrupts, such that page access patterns can be deterministically monitored for every enclave instruction, giving precise insights into the operation of a victim enclave.
+By conservatively under-estimating the APIC timer interval, SGX-Step results in either zero or single-steps, but always avoids multi-steps [36, Table 5].
+Subsequent works [37, 38, 39, 40, 41] showed that the accessed (A) bit in the enclave code page deterministically distinguishes single-steps (enclave instruction retired; A=1) from zero-steps (A=0).
+This ability to perfectly and deterministically single-step production enclaves using SGX-Step prompted Intel to develop the opt-in AEX-Notify hardware-software mitigation (cf.
+Section˜VII).
+To date, instruction-granular page-access traces extracted with SGX-Step have been repeatedly abused to reliably exploit enclave interface vulnerabilities [40] or to reconstruct cryptographic key material [38, 39].
+In the following, we show that such traces are sufficient to recover weights from real-world NN libraries running inside a production SGX enclave.
+III System and Adversary Model
+Attacker Capabilities.
 We adhere to Intel’s standard SGX threat model, where the adversary has full root access on the victim machine, as is e.g., the case when deploying an ML model on external cloud infrastructure.
-Notably, we do not require physical access to the machine for our attack.
-Privileged software adversaries can use SGX-Step (cf.
-Section 2) to precisely single-step a target enclave and obtain instruction-granular page access traces.
-We discuss generalizations to other TEEs in Section 8
-We assume the network’s weights and biases are confidential (sealed), while the enclave source code is known.
-The latter is a common assumption in SGX attacks [xu_controlled-channel_2015, van_bulck_sgx-step_2017, VanBulck17, Brasser17, shinde2016preventing] and the default case in the Intel SGX architecture and SDK.
-Non-standard confidential-code deployments only add “security through obscurity” by requiring an additional, out-of-scope reversing phase [puddu2024lack, yan_cache_2020].
-We assume a pre-trained ML network has been securely loaded into an SGX enclave to perform inference.
-We do not consider the security of the model training process, which is orthogonal to our attacks, but believe it to be an interesting avenue for future research.
+The attacker is a software-only adversary without local physical access, for instance they may gain root access remotely over the network.
+Such privileged software attackers can send arbitrary inputs through the enclave software interface and can further interrupt the enclave by manipulating page tables and interrupts (cf.
+Section˜II).
+For the latter, we leverage the widely-used SGX-Step [6] framework to precisely single-step a target enclave and obtain instruction-granular page access traces.
+Notably, while attackers may debug enclaves using a copy of the victim code to facilitate attack development, we conduct all final attacks with SGX-Step on target enclaves running in SGX production mode, i.e., without access to debug features.
+While our work focuses on the Intel SGX architecture, the underlying insights and weight-recovery techniques are applicable to other TEEs.
+Notably, the required side-channel primitives offered by SGX-Step, i.e., precise single-stepping and page accesses, have been demonstrated on alternative VM-based TEEs like AMD SEV [42] and Intel TDX [43, 44].
+Target Model.
+We assume a pre-trained ML network.
+We do not consider the training process, which is orthogonal to our attack (but we believe it to be an interesting avenue for future research).
+We focus on Feedforward Neural Networks because they effectively illustrate the problem we exploit and are easily deployable in SGX enclaves (e.g., without dynamic linking or requiring libc) using Tensorflow Microlite.
+Lastly, we assume the network architecture (at least the number of neurons per layer and activation functions) is known, e.g., it is an open-source or well-known architecture or a transfer-learning model, or can be recovered via a cache attack [15].
+Target Enclave.
+We assume that the network has been securely loaded into a production-mode SGX enclave to perform “secure” inference.
+The network’s weights and biases are confidential (sealed), while the enclave source code is known.
+The latter is a common assumption in SGX attacks [10, 6, 35, 33, 34] and the default case in the Intel SGX architecture and SDK.
+Non-standard confidential-code deployments only add “security through obscurity” by requiring an additional, out-of-scope reversing phase [45, 15].
+Since the attacker never has access to a self-contained binary that contains the hidden parameters, an attack such as the one proposed by [46] is not possible [46].
 Inputs are passed into the enclave and the output is returned outside the enclave through the standard ecall interface.
-Based on the security guarantees of SGX, an ML developer would assume that their network is protected against duplication, as the inference is handled entirely within an enclave, and weights are never stored outside.
-We focus on Feedforward Neural Networks because they effectively illustrate the problem we exploit and are easily deployable in SGX enclaves (e.g., without dynamic linking or requiring libc) using Tensorflow Microlite.
-We also assume the network architecture (at least the number of neurons per layer and activation functions) is evident from the source code, known (as in the case of a well-known architecture or a transfer-learning model), or can be recovered via a cache attack [yan_cache_2020].
-4 Vulnerable Activation Functions
-In this paper, we mostly consider activation functions that include a call or calls to a functions in the exp() family, e.g., sigmoid(), tanh(), and expf().
-This is based on our initial observation that typical implementations of exp() contain a number of input-dependent branches with different return cases.
-To give an example, LABEL:lst:expf_in_tlibc shows some different code paths of the implementation of expf() in SGX’s tlibc [^1] https://github.com/intel/linux-sgx/blob/7385e1/sdk/tlibc/math/e_expf.c.
-1float expf(float x)
-2{
-3 // ...
+Based on the security guarantees of SGX, an ML developer would assume that their network is protected against duplication, as inference is handled entirely within an enclave and weights are never stored outside.
+Figure 2: Visualization of expf() CPU instruction count regions by input.
+The Underflow and Overflow regions have different instruction counts: 18 and 17 respectively.
+Note also the symmetrical regions of Inner1, inner2, and Inner3.
+Activation Functions.
+Finally, our attack requires that the victim network use a vulnerable activation function which contains at least two input-dependent branches with unique step counts such as the time they take to return.
+To give an example, LABEL:lst:expf_in_tlibc shows some different code paths of the implementation of expf() in SGX’s tlibc [47].
+Observing the memory access/execution patterns, e.g., with SGX-Step, allows an attacker to learn which return case an input results in, and thus obtain information on the input to exp().
+For this reason, we mostly consider activation functions that include a call or calls to a functions in the exp() family, e.g., sigmoid(), tanh(), and softmax().
+Notably, such activation functions are widely used in transformer architectures (even though we do not consider attacks on transformers in this paper), as with sigmoid() in Deepseek V3 [48], GELU() in BERT [49] and SwiGLU() [50] in Llama [51].
+softmax() was specified in the initial transformer paper [52] and is still used in many of these implementations.
+Note that while several variants of exp() implementations exist, they commonly feature the same structure, where multiple error cases are handled with early return statements.
 1 if(hx >= 0x42b17218) { /* if x>=88.721... */
@@
 5 if(x < u_threshold) return ...; // underflow
-9 }
 6 } if(hx > 0x3eb17218) {
-11 if(hx < 0x3F851592) {
-12 // ...
-13 } else {
-14 // ...
-15 }
-16 }
+7 if(hx < 0x3F851592) { /* ... */ } else { /* ... */ }
 8 } else if(hx < 0x31800000) { /* when x<2**-28 */
@@
 10 }
-20 // ...
-21}
 Listing 1: Excerpt from expf() in tlibc.
-Observing the memory access/execution patterns, e.g., with SGX-Step, allows an attacker to learn which return case an input results in, and thus obtain information on the input to exp().
-Note that while several variants of exp() implementations exist (both in terms of algorithms for calculation and in terms of single and double precision floats), many of them feature the same structure, where several cases are handled with different return statements.
 Attacking an activation function containing a call to exp() is similar to attacking exp() directly.
-One must be careful to note the sign of the input (as in the case of the $-x$ in sigmoid()’s $\exp\left(-x\right)$) and search in the correct direction.
-In the case of functions like tanh(), which include multiple discrete calls to exp(), the code to process the memory access traces can not assume a one-to-one relationship between a call to exp() and a single neuron’s activation.
-Another important consideration is the expressive power of the function, which we discuss in more detail in Section 5.3.
-We summarise common activation functions in Table I, and discuss their security in Section 7.2.
-Despite our focus on the exp() family of functions, any function exhibiting measurable input-dependent variation in instruction counts (or memory access patterns) is likely vulnerable.
-As long as the attacker can force a vulnerable function to process inputs that fall into multiple distinguishable classes, it is possible to search for the boundary points between these cases, as we discuss in detail in Section 5.
-In particular, relu() does not rely on a call to exp(), but to conditional logic or a call to std::max().
-While relu() is ubiquitous, especially in ML-aided computer vision applications, transformers and other modern Large Language Models use a variety of exp()-based activation functions such as softmax(), tanh() and GELU() in addition to relu().
-We discuss the vulnerability of relu() in Section 7.2.1.
+One must be careful to note the sign of the input (as in the case of the x-x in sigmoid()’s $\exp\left(-x\right)$) and search in the correct direction.
+In the case of functions like tanh(), which can include multiple discrete calls to exp(), the code to process the memory access traces cannot assume a one-to-one relationship between a call to exp() and a single neuron’s activation.
+Another important consideration is the expressive power of the function, which we discuss in more detail in Section˜IV-E.
+We note NNs may access the maths functions that underpin their activations from a few different sources.
+Whether defined by an ML framework itself or imported externally from low level standard libraries such as glibc’s libm or even dedicated high performance maths libraries like Eigen (which may be vectorised to exploit hardware acceleration features), the code may be vulnerable to our attack.
+We explore this with our second case study in Section˜V-B and will more generally consider how the source of these functions can affect their vulnerability to our attack in Section˜VI.
+Finally, despite our focus on the exp() family of functions, any function exhibiting measurable input-dependent variation in instruction counts (or memory access patterns) is likely vulnerable.
+As long as the attacker can force a vulnerable function to process inputs that fall into multiple distinguishable classes, it is possible to search for the boundary points between these cases, as we discuss in detail in Section˜IV.
+In particular, relu() does not rely on a call to exp(), but on conditional logic or a call to std::max().
+We summarise common activation functions in Table˜I and specifically discuss the vulnerability of relu() in greater detail in Section˜VI.
+IV Methodology
+First, we describe the overall concept for our attack (Section˜IV-A) against a NN and then consider (Section˜IV-B) different ways ML frameworks may implement activation functions.
+Next (Section˜IV-C), we show how to perform our attack practically using SGX-Step to generate per-neuron traces.
+We then explain in more detail how to attack the first layer (Section˜IV-D) and the subsequent layers (Section˜IV-E).
+Finally (Section˜IV-F) we estimate the performance of our attack.
+IV-A Attack Concept
+The starting point for our attack is the above observation (cf.
+Figure˜2) that, unlike in cryptographic libraries, the maths functions that underpin many ML activation functions have not been hardened to ensure they execute in constant time and without input-dependent branches and memory access patterns.
+A classic example in the cache timing literature is a Look-Up Tables that may be accessed in deterministic patterns based on the function inputs.
+We believe this work has not been undertaken in the context of ML due to performance concerns or because security is seen as less of a focus.
+By examining the underlying functions statically, dynamically, or by reading the code (if available), we can deduce the possible branches through the function and the timing as well as memory access patterns therein.
+Also of interest are cases where certain input values are handled differently (via multiple return statements).
+A special case of those are early returns, where certain inputs cause the function to return before the full output is computed.
 TABLE I: Survey of Potentially Vulnerable Activation Functions.
@@
 sigmoid() $(0,1)$
-tanh() $(-1,1)$
+tanh() (1,1)(-1,1)
 softplus() $(0,\infty)$
@@
 relu() $(0,\infty)$
-4.1 ML Framework Functions
-Thus far, our discussion of the functions used for activation functions have focused on low level mathematical used by ML framework developers inside of higher level framework code.
-A Tensorflow neuron with tanh() activation might not, for instance, directly return the output of a call to math library’s implementation of tanh(); there might be wrapping or processing of the result.
-This is of interest as we can also exploit input-dependent behavior at this higher framework or application level.
-An example of this is the code to Tensorflow lite’s sigmoid() function (called Logistic() in the source), shown in LABEL:lst:sigmoid_in_tflite in Appendix B.
-We can see that on top of any side-channel leaks due to calls to std:exp(), a higher-level set of memory access patterns and instruction count cases are introduced by a conditional routing to different functions (or simply a return of 1) based on the input.
-In this way, this sigmoid() function can actually be attacked at two levels which we will describe in Section 6.2.
-5 Methodology
-First, we describe the overall concept for our attack (Section 5.1) against a neural network.
-We then explain in more detail how to attack the first layer (Section 5.2) and finally the subsequent layers (Section 5.3).
-5.1 Attack Concept
-The starting point for our attack is the above observation (cf.
-Section 4) that, unlike in cryptographic libraries, the math functions that underpin many ML activation functions have not been hardened to ensure they execute in constant time and without input-dependent branches and memory access patterns [^2] A classic example in the cache timing literature is Look-Up Tables that may be accessed in deterministic patterns based on the function inputs..
-We believe this work has not been undertaken in the context of ML due to performance concerns or because security is seen as less of a focus.
-By examining the underlying functions statically, dynamically, or by reading the code (if available), we can deduce the possible branches through the function and the timing as well as memory access patterns therein.
-Also of interest are cases where certain input values are handled differently (via multiple return statements).
-A special case of those are early returns, where certain inputs cause the function to return before the full output is computed.
-$i_{1}$$w_{1}$$i_{2}$$w_{2}$$\Sigma$$f(\Sigma)$ Activation function $output$ Bias $b$
-Figure 2: Visualization of a single 2-input neuron. We use $\Sigma$ in this paper to represent the sum of all $i_{i}*w_{i}$ and $b$ passed into an activation function.
-Given that such code patterns take different numbers of instructions (and exhibit different memory access patterns) depending on the input, we can more generally assert that there is a correlation between different code paths and the execution trace.
+Given that such code patterns take different numbers of instructions (and exhibit different memory access patterns) depending on the input, more generally there is a correlation between different code paths and the execution trace.
 Though these differences might be subtle (e.g., a single instruction), they are measurable given a suitable side channel.
 Thus, if an attacker can determine which branch a function has taken, they learn something about the input or argument to that function.
-As mentioned before, one function with these properties is expf().
-Figure 3 shows the 11 different return cases of this function based on input and their associated CPU instruction counts (collected using gdb).
-We note that we found other optimised implementations of expf() that all use a similar algorithm and thus exhibit the same input-dependent memory access patterns.
-We focus here on the one used by SGX tlibc.
-For ease of discussion we labeled these cases using the names shown.
-For example, expf(4) puts us in case Normal (+)(+)( + ), whereas expf(110) results in the Overflow case.
-We define a threshold (shown in Figure 3 as dotted black lines) as an input value that is on the precise border of two output classes.
-For example, $88.72168$ is the value between Normal (+)(+)( + ) and Overflow.
-We can also use knowledge of these regions and thresholds in the other direction: for example, if we know that an execution of expf() takes 17 CPU instructions (placing us somewhere in the Overflow region), we know the input to that function call must be greater than $88.72168$.
+As stated, one function with these properties is expf().
+Figure˜2 shows the 11 different return cases of this function based on input and their associated CPU instruction counts (collected using gdb).
+We focus here on the expf() implementation from SGX’s tlibc, but we note that we found other optimised implementations all using a similar algorithm and thus exhibiting similar input-dependent memory access patterns.
+For ease of discussion we have labeled these 11 cases using the names shown in Figure˜2.
+For example, expf(4) results in case Normal $(+)$, whereas expf(110) results in Overflow.
+We define a threshold (shown in Figure˜2 as dotted black lines) as an input value that is on the precise border of two timing classes.
+For example, the value $88.723$ is between Normal $(+)$ and Overflow.
+We can also use knowledge of these regions and thresholds in the other direction: for example, if we know that an execution of expf() takes 17 instructions (in the Overflow region), we know the input to that function call must be greater than $88.723$.
 We note that the cases for Overflow and Underflow are both early returns and as such their instruction counts are much lower than the other ‘normal’ returns which actually perform some sort of unique computation.
@@
 If we consider the scenario where an input to expf() is being multiplied by a hidden constant $c$ and we have access to the instruction count of the execution, we can reveal $c$ by finding (via e.g., a binary search) a threshold value between two instruction count classes and then dividing that value by our input.
-For example, if we find that the input $3.2164$ multiplied by $c$ leads to threshold value $88.72168$, $c$ must be $\nicefrac{{88.72168}}{{3.2164}}\approx 27.5842$.
+For example, if we find that the input $3.2164$ multiplied by $c$ leads to threshold value $88.723$, $c$ must be $\nicefrac{{88.723}}{{3.2164}}\approx 27.5846$.
 We therefore define $3.2164$ as a convergence point: an input to the system that causes a threshold value and thus a leak of the precise input to the function.
-Figure 3: Visualization of expf() CPU instruction count regions by input.
-The Underflow and Overflow regions have different instruction counts: 18 and 17 respectively.
-Note also the symmetrical regions of Inner1, inner2, and Inner3.
-We now extend the above into a chosen input attack against a toy ML network.
-Consider an attack against a simple neural network with only a single neuron, as shown in Figure 2.
+[…]
+Figure 3: Visualization of a single 2-input neuron. We use $\Sigma$ in this paper to represent the sum of all $i_{i}*w_{i}$ and $b$ passed into an activation function.
+We now extend the above into an attack against a toy ML network: consider a simple NN with only a single neuron, as shown in Figure˜3.
 The activation function is simply expf().
@@
 We are free to choose values for $i_{1}$ and $i_{2}$.
-We first note that the input to a neuron’s activation function is the sum of the products of various user-controlled inputs and their hidden weights, plus a hidden bias term ($\Sigma$ in Figure 2); this is similar to the case described previously except we have two hidden constants multiplied by two different inputs and an additional term, the bias, that is never multiplied with our inputs.
+We first note that the input to a neuron’s activation function is the sum of the products of various user-controlled inputs and their hidden weights, plus a hidden bias term ($\Sigma$ in Figure˜3); this is similar to the case described previously except we have two hidden constants multiplied by two different inputs and an additional term, the bias, that is never multiplied with our inputs.
 If there was no bias term, we could simply use a variant of the approach described above and iterate through each input (setting all other inputs to zero, allowing us to focus on one weight at a time) and recovering each parameter independently.
 The addition of the bias term adds another unknown which means we can only find through constraints on the value of the hidden weights and bias:
-$i_{1}*w_{1}+0*w_{2}+b=88.72168$
-$\Rightarrow w_{1}=\frac{88.72168-b}{i_{1}}$
+[…]
 To solve for all three unknowns, we find enough sets of values (one for each input to the neuron) to generate a system of equations to solve for as many unknowns as we have (number of weights plus bias term), for example:
-[…]
+$i_{1}*w_{1}+\cancel{0*w_{2}}+b$ $=threshold_{1}$
 $\cancel{0*w_{1}}+i_{2}*w_{2}+b$ $=threshold_{2}$
 $i_{3}*w_{1}+i_{4}*w_{2}+b$ $=threshold_{3}$
-[…]
-[…]
-Finally, if we replace exp() above with a different function (such as sigmoid()) that also leaks threshold values, the only thing that changes is how to conduct the binary search: the recovery procedure is otherwise identical.
+Here, we define $(i_{1},\,0)$, $(0,\,i_{2})$, and $(i_{3},\,i_{4})$ as convergence point sets, because each set of chosen inputs forces the activation function to a threshold value.
+Once we have enough equations (one for each unknown), we solve using linear algebra for the unknown weights and biases.
+Note that the thresholds do not have to be different, though the more thresholds are hit, the more robust the solution.
+Though it is possible to find convergence point sets like $(i_{1},\,0)$ and $(0,\,i_{2})$ above where all but one input is locked to zero and the remaining input is used to search for a threshold point, this is not required.
+In fact, by locking all but one input to some small magnitude random numbers (e.g., $i_{3}$ in the third equation) and searching on the single unlocked parameter ($i_{4}$), an arbitrary number of convergence point sets can be found.
+Finally, if we replace exp() above with a different function (such as sigmoid()) that also leaks threshold values, the only change is how to conduct the binary search; the recovery procedure is otherwise identical.
 Building on this concept, we will show how it is possible to recover the hidden parameters of more complex networks to a high degree of accuracy.
-5.2 Full Attack on First Layer
-Continuing on from Section 5.1, we discuss how to attack the first layer of a production neural network.
+IV-B ML Framework Functions
+Thus far we have largely assumed a low level mathematical definition for these functions.
+However, as discussed in Section˜III, a Tensorflow neuron with a tanh() activation might not directly return the output of a call to the underlying maths library’s implementation of tanh(); there might be wrapping or processing of the result at the framework level.
+This is of interest as we can also exploit input-dependent behaviour at this higher framework or application level.
+An example of this is Tensorflow lite’s sigmoid() function (called Logistic() in the source), shown in Appendix˜A.
+We can see that on top of any side-channel leaks due to calls to std:exp(), a higher-level set of memory access patterns and instruction count cases are introduced by a conditional routing to different functions (or simply a return of 0) based on the input.
+In this way, this sigmoid() function can actually be attacked at two levels which we will describe in Section˜V-B.
+IV-C Capturing Traces with SGX-Step
+To practically execute our attack, we used SGX-Step to collect the CPU instruction counts necessary to understand the state of our victim networks and apply the steps described above.
+This required that our networks run inside an SGX enclave.
+Given the constraints of SGX’s standard library replacement tlibc, we decided to employ Tensorflow Microlite, a software package for running (converted) Tensorflow lite models on microcontrollers, instead of “full” Tensorflow (which relies on a variety of system calls not easily provided by SGX).
+Our SGX enclave thus included a statically compiled Tensorflow Microlite library to perform inference.
+We converted our tensorflow models to Tensorflow lite and then further exported them as to a Tensorflow Microlite byte array which can be compiled directly into a C program.
+These byte arrays are added to the enclave code which allows us to perform “secure” inference inside SGX.
+We note that for practical deployments of SGX-protected ML, using such a reduced variant of Tensorflow would ease deployment, reducing code and thus Trusted Computing Base (TCB) size, while maintaining compatibility with Tensorflow-trained models.
+The underlying implementation [47] of exp() used by our network during inference comes from SGX’s libc alternative, tlibc.
+We first profiled the execution of Tensorflow Microlite on an attacker-controlled debug enclave, so that we could understand execution flow by directly observing call traces via the interactive sgx-gdb debugger.
+An example can be seen in Appendix˜B.
+This provides an intuition for how Tensorflow Microlite performs inference and which functions and pages are accessed.
+Figure 4: Inset of an SGX-Step instruction-granular page access trace showing the execution of the first layer of a NN with 16 neurons.
+Note that the intra-page instruction pointer (RIP) values (recovered by putting an enclave in debug mode) are displayed here only for reference purposes and are not used in our attack.
+We include them to show that knowing the current page provides enough information to interpret traces without precise RIP values.
+We then developed an attack application, which starts and interacts with the victim enclave in production mode, i.e., without access to debug features.
+We use this to record page accesses which we later interpret into activation function states.
+Using SGX-Step, we log all page accesses within the production enclave (after a certain trigger page was accessed) and interpret the results as part of a post-processing step.
+Unlike sgx-gdb, SGX-Step only offers a relatively coarse-grained spatial resolution of 4 KiB memory pages, which prevents production enclave attackers from directly identifying the specific symbol (e.g., function or data) accessed by the enclave code.
+However, SGX-Step’s precise instruction-level temporal resolution enables the annotation of coarse-grained page-access traces with the exact number of instructions executed on each page.
+As illustrated in Figure˜4, this capability allows to accurately reconstruct the inference execution structure, even without knowledge of the lower 12-bit page offset of the secret enclave instruction pointer (ERIP).
+By comparing the dotted red line which depicts the true ERIP values available to sgx-gdb and the black line which shows the reduced page-level granularity, one can see how little is lost.
+It is worth nothing that each page can contain multiple functions and so it is difficult to assign a precise “meaning” to each page without context.
+The instructions executed on one page could pertain to two very different activities at different points in execution.
+However, the combination of coarse-grained page-access patterns along with the exact amount of instructions executed per page, annotated in black text, clearly suffices to identify individual function calls or data accesses.
+In practice, this means that even if multiple symbols are co-located on the same 4 KiB page, we can still interpret the logs.
+We also know precisely the length (or length range) of different runs at different points during the execution, i.e., we can also tell if a trace is invalid and immediately reject it.
+By studying inference traces from the perspective of both sgx-gdb and SGX-Step (raw traces or visualizations like those shown in Figure˜4), clear patterns emerge.
+For example, the function ExpEval() is located on a different 4 KiB page than expf() (each 0x1000 bytes is a new page) and we can plainly interpret the movement between them as the higher level neuron code invoking the lower level expf() maths function 16 times (once for each neuron in the first layer of the model this plot was constructed from).
+The 18 instruction count long runs at the extreme left and right on page 0x2c000 seem to correspond to the start and end of a layer, which makes parsing each layer trivial.
+We note that the step counts provided by sgx-gdb and SGX-Step may vary slightly (but deterministically) due to the documented effects of macro-op fusion [38].
+The final part of our tooling is a Python script and library which further abstracts and manages invocations of the enclave program.
+Because each neuron is accessed in order and each return occurs in sequence, it should be possible to monitor each neuron’s step count independently.
+In order to orchestrate the full attack, we use this script to interpret the log output by SGX-Step into a per-neuron state.
+Even though we receive the state for every neuron in each log, we only focus on one at a time.
+It takes about ten seconds to run each input through our SGX-Step setup and parse the trace.
+An example output of the running attack is shown in LABEL:lst:neuron_centric_attack in Appendix˜C.
+IV-D Neuron-Centric Attack on First Layer
+Here we discuss how to convert the exploit discussed in Section˜IV-A into a full attack on the first layer of a production NN.
 As before, the goal of this phase is to find a number of convergence point sets equal to the number of inputs plus one in order to have a system of equations we can solve to recover the weights and bias.
@@
 If there is filtering or normalization, however, we treat the first layer like all subsequent layers.
-If we can inject arbitrary values into the first layer, we iterate over every neuron in the first layer and recovery its hidden parameters in 2 phases: calibration and binary search.
-5.2.1 The Calibration Phase
-In the first phase, we quickly produce sensible starting points and context for the second phase of the attack.
-We prepare a test input to the network that is all small non-zero values but set the first value of the array to a large value.
-If we run this test input and detect that the neuron overflowed, we know that this input is large enough to cause an overflow and that the weight between the first input and this neuron must be positive.
-If instead we produce an underflow, we know that the weight must be negative.
-If we are not in an overflow or underflow state (perhaps owing to a low magnitude weight).
-In this case, we simply multiply the large value in our test input by 10 and try again.
+If we can inject arbitrary values into the first layer, we can conduct a chosen-input, neuron-centric attack pass by iterating over every neuron in the first layer and recovering their hidden parameters in 2 phases: calibration and binary search.
+IV-D1 The Calibration Phase
+Because the attack is a binary search to find hidden parameters of unknown value, can achieve dramatic speedups if we can find a reasonable upper bound to start the search from, rather than stepping down from the maximum possible float value.
+It is also useful to know the signs of the weights connected to our neurons.
+We can recover both pieces of information through the calibration step.
+We prepare a test input to the network that is all small, non-zero values except for the first value which is a very large value.
+Next, we perform inference on this test input and consult the trace to see if the target neuron overflowed or underflowed.
+If it has, we know both that this input is already large enough to cause an overflow/underflow and the sign of the weight between the first input and this neuron.
+If the neuron has not overflowed or underflowed, (owing to a comparatively lower magnitude weight) we simply multiply the large value in our test input by 10 and try again.
 We continue in this way until we find the sign and magnitude required to over/underflow this neuron from each input to the network.
 Note that we do not actually binary-search towards a threshold between two cases here; we merely attempt to reach an overflow/underflow case.
-In this case, we do not need to find a threshold for the instruction count leak to be informative.
+It is not necessary to find precise threshold points to learn the rough magnitude and sign of the inputs.
 Recovering the signs for each weight is useful since some of the cases (most of the “normal” ones) are symmetrical about zero, meaning we cannot learn the sign of the input only based on the state.
 Recovering the sign before we begin the binary search means we always know exactly where we are in the state space.
-Recovering a value that overflows/underflows the neuron is useful to accelerate the attack.
-Though we could choose the maximum float value for the starting point of our binary search, the lower the initial guess, the fewer steps the search has to take.
-Searching "up" in the calibration step and then searching down in the second phase reduces the number of invocations needed.
-5.2.2 The Binary Search Phase
-Armed with the information collected in the calibration phase, we continue the attack against the same neuron in phase 1.
-Now, we can binary search to find the desired number of convergence point sets for this neuron.
+Searching “up” in this calibration step gives us a sensible starting point to search down and thus reduces the number of invocations needed in the next phase.
+IV-D2 The Binary Search Phase
+Armed with the information collected in the calibration phase, we continue the attack against the same neuron as in phase 1 by finding the desired number of convergence point sets.
 For as many convergence sets as we need, we start by generating an input set that is made of $n$ locked random values (one for each input) except that the $i$’th input is instead set to the large positive value recovered during calibration.
-We feed the input into the network and note the return case by interpreting the CPU instruction counts for each neuron.
+We feed the input into the network and note the return case by interpreting the CPU instruction counts trace for the neuron.
 Even if some of the weights have different signs, the overall $\Sigma$ should be dominated by the large input multiplied by its weight.
@@
 If the neuron has instead underflowed, we must be in the boundary between underflow and normal (-).
-If the input is negated before entering exp() as in sigmoid(), the direction to search should be reversed.
-We continue the binary search on that $i$’th input (searching back up if we fall below an underflow/overflow case in terms of magnitude) until we hit the desired search depth or our search does not change, because we reached the limit of decimal places for a floating point number.
-The tradeoff between depth and speed is shown in Table II.
+If the activation’s input is negated before being passed into exp() (as in sigmoid() ), the direction to search should be reversed.
+We continue the binary search on that $i$’th input, searching back up if we fall below an underflow/overflow case in terms of magnitude.
+Eventually the search ends either because the desired search depth was reached or because we reached the limit of precision for the underlying floating point maths.
 Using this algorithm, we can generate as many novel convergence sets as we require since the unchanging members of the input set are all randomised between inputs.
@@
 Note however that it is not required to target the thresholds between underflow/normal and normal/overflow; these are just the easiest to find since we can force an under/overflow and then work our way up/down until we find the normal region.
-There are other thresholds within exp() (as shown in Figure 3) that can be exploited if they can be reached so long as we know the sign of the weight and thus in which precise state we are in.
-Once we have the requisite number of equations, we can solve the system of equations as described in Section 5.1 to recover the weights/biases for this first neuron.
+There are other thresholds within exp() (as shown in Figure˜2) that can be exploited if they can be reached so long as we know the sign of the weight and thus in which precise state we are in.
+Once we have the requisite number of equations, we can solve the system of equations as described in Section˜IV-A to recover the weights/bias for this first neuron.
 We then perform the exact same two-part algorithm on each remaining neuron until we solve the entire layer.
-The output from our recovery tool as it recovers one neuron is depicted in Appendix A.
-5.3 Extending the Attack to Deeper Layers
+The output from our recovery tool as it recovers one neuron is depicted in Appendix˜C.
+IV-E Extending the Attack to Deeper Layers
 In order to extend the attack into subsequent layers, we must solve two key problems.
-First, how can we ‘unwrap’ the effect of already-solved early layers to insert arbitrary inputs beyond them, deeper into the network?
+First, how can we mitigate or ‘unwrap’ already-solved early layers to insert arbitrary inputs beyond them, deeper into the network?
 And second, is it always possible to generate those values?
-Unwrapping Solved Layers
+Unwrapping Solved Layers.
 Our attack methodology for the first layer relies on being able to precisely craft a set of “target” in order to search towards threshold points.
 In order to be able to do this for deeper layers, we first have to “unwrap” all the recovered layers between the input layer and the layer we target.
-In the following set of equations, we show how to rearrange the normal operation of a neuron to isolate the input as opposed to the target.
-$act(input*W+b)$ $=target$
-$input*W+b$ $=act^{-1}(target)$
-$input*W$ $=act^{-1}(target)-b$
-[…]
-$input$ $=(act^{-1}(target)-b)*W^{-1}$
-Note that it is also possible to do this without matrix inversion for greater numerical efficiency, e.g., using numpy.linalg.lstsq(W, act_inv(target)-b).
-In these equations, $act$ and $actinv$ are known since we know the architecture of the network and $W$ and $b$ are known because we already solved this layer.
-Therefore, given a solved layer’s hidden parameters, a desired ‘target’ set of values, and the appropriate inverse activation function, we can calculate the input we must feed into the layer to produce a desired target as the output of that layer.
-Note also that the above assumes the activation function is invertible without a substantial loss of accuracy.
-Inverting these functions might lead to an inability to pass certain values forward, which we discuss below.
-Expressive Power
-Even if we can algebraically calculate how to pass arbitrary values through a solved layer, subsequent layers are harder to attack because we are limited by the “expressive power” (see Table I) of the neurons before the target layer in terms of both magnitude and sign.
+In Appendix˜D, we show the simple linear algebra required to rearrange the normal operation of a neuron to isolate the input as opposed to the target.
+In these equations, actact and actinvactinv are known since we know the architecture of the network and $W$ and $b$ are known because they belong to a solved layer.
+Therefore, given these recovered parameters, a desired ‘target’ set of values, and the appropriate inverse activation function, we can calculate the input we must feed into the layer to produce a desired target as the output of that layer.
+Note this assumes the activation function is invertible without a substantial loss of accuracy; inverting these functions might lead to an inability to pass certain values forward, which we discuss below.
+Expressive Power.
+Even if we can algebraically calculate how to pass arbitrary values through a solved layer, subsequent layers are harder to attack because we are limited by the “expressive power” (see Table˜I) of the neurons before the target layer in terms of both magnitude and sign.
 There are two consequences to this.
@@
 The second is that certain thresholds such as those between expf()’s overflow/underflow and normal cases have a high magnitude ($\approx 88.723{-}103.972$) relative to generally smaller magnitude weights and biases.
-sigmoid() has a very small output range ($0{-}1$) making it impossible to produce a large enough $\Sigma$ in the next layer to ‘reach’ one of these two thresholds unless either the network is very dense with small-magnitude positively weighted neurons or has some positive and large magnitude ($>1$) weights connected to it.
+sigmoid() has a very small output range (010{-}1) making it impossible to produce a large enough $\Sigma$ in the next layer to ‘reach’ one of these two thresholds unless either the network is very dense with small-magnitude positively weighted neurons or has some positive and large magnitude (>1>1) weights connected to it.
 This suggests denser networks are more expressive and therefore more vulnerable.
-In both of these cases, we rely on certain key hidden parameters, which means that certain networks might not be vulnerable to our attack.
-We however note that security guarantees should never assume a certain network architecture.
-Attack Outline
+In both of these cases, we rely on certain network properties.
+Though this means that certain networks might not be vulnerable to this method, security guarantees should never rest on architectural decisions.
+Attack Outline.
 Nevertheless, extending the attack into deeper layers is possible with certain caveats.
-There is no need to focus only on the high magnitude thresholds we previously discussed between the underflow/overflow cases and the normal one; there are six in the range from $\approx$ -1.04–1.04 (see Figure 3).
-By searching anywhere within this region (e.g., near the subnormal case as in [Gongye20]) we can, as before, build convergence sets for our inputs and solve for the hidden parameters.
-Because we might pass through a low expressiveness activation function like sigmoid (output range from 0–1), which is likely being further reduced by multiplication with a weight with magnitude <1, we might not cross a single threshold point (the presence of the bias as a component of $\Sigma$ means that we might be shifted away from zero far enough that we cannot reach both sides of the subnormal case).
+There is no need to focus only on the high magnitude thresholds we previously discussed between the underflow/overflow cases and the normal one; there are six in the range from $\approx$ -1.04–1.04 (see Figure˜2).
+By searching anywhere within this region (e.g., near the subnormal case as in [27]) we can, as before, build convergence sets for our inputs and solve for the hidden parameters.
+Because we might pass through a low expressiveness activation function like sigmoid() (output range from 0–1), which is likely being further reduced by multiplication with a weight with magnitude <1, we might not cross a single threshold point (the presence of the bias as a component of $\Sigma$ means that we might be shifted away from zero far enough that we cannot reach both sides of the subnormal case).
 As such, we instead start with a grid search to scan (at some resolution up to the attacker which trades off speed and the chances of finding a narrow instruction count class) for multiple instruction count classes.
@@
 If we are able to reach multiple classes, we can then perform a binary search between those classes to isolate the threshold and build a convergence point set, as before.
-6 Implementation and Orchestration
-We trained several neural networks using “full” Tensorflow (in Python).
-In this section, we will first introduce a proof-of-concept attack on a case-study victim network with an expf() activation function.
-Next, we show how to attack a more realistic activation function and start to recover partial information from deeper layers.
-6.1 Regression Model with expf()
-Model
-We first focus on a regression model to predict home insurance costs, which was trained on the “Medical Cost Personal Dataset” [choi18].
+IV-F Estimating Attack Performance
+We note that the number of executions to perform our attack is dependent on the architecture of the network (and therefore the number of parameters to recover) but also some factors the attacker can control such as the depth of the binary search.
+Naively, the total number $Q$ of queries to recover the first layer can be computed as: $Q=\,P\,\cdot\,N\,\cdot S$, where $P$ is the number of hidden parameters (in the first layer, e.g., one weight per input plus one for the bias) to solve for, $N$ is the number of neurons in the first layer, and $S$ is the maximum number of steps in the binary search, minimised by the calibration process described in Section˜IV-D1.
+The $P$ value correlates to the minimum number of equations we require to solve a system of equations: one for each unknown.
+We note that we can compensate for noise or other errors by adding extra equations to produce better solutions at the cost of more queries.
+In testing, finding 3 additional convergence sets helped reduce the maximum error by 28% at depth 25 compared to an attack where precisely $P$ convergence sets were generated (this is visualised in Appendix˜K).
+It might be possible with future study to detect particularly weak systems of equations analytically and dynamically add extra convergence sets to improve the results.
+V Experimental Evaluation
+Here we introduce and share the results of our attack against three PoC networks trained on different datasets in “full” Tensorflow (in Python) and then converted to Tensorflow Microlite using the process described in Section˜IV-C.
+V-A Minsurance: A Regression Model using expf()
+Model.
+We first focus on a regression model to predict home insurance costs, which was trained on the “Medical Cost Personal Dataset” [53].
 The network has three hidden layers.
@@
 A final layer contains a single relu() neuron.
-For this first network, we only focus on the first hidden layer. Despite that, we note that:
+Recovery. For this first network, we only focus on the first hidden layer. Despite that, we note that:
   1. 1.
@@
     The attack is agnostic to the accuracy of the network.
-Victim Enclave
-The enclave code runs inside of SGX.
-Given the constraints of SGX’s standard library replacement tlibc, we decided to employ Tensorflow Microlite, a software package for running (converted) Tensorflow lite models on microcontrollers, instead of “full” Tensorflow (which relies on a variety of system calls not easily provided by SGX).
-Our SGX enclave thus included a statically compiled Tensorflow Microlite library to perform inference.
-We converted our tensorflow models to tflite and then further exported them as to a byte array which can be compiled directly into a C program.
-These byte arrays are added to the enclave code which allows us to perform “secure” inference inside SGX.
-We note that for practical deployments of SGX-protected ML, using such a reduced variant of Tensorflow would ease deployment, reducing code and thus Trusted Computing Base (TCB) size, while maintaining compatibility with Tensorflow-trained models.
-The underlying implementation [sun_microsystems_tlibcs_1993] of exp() used by our network during inference comes from SGX’s libc alternative, tlibc.
-Attack Code
-To practically execute our attack, we used SGX-Step to collect the CPU instruction counts necessary to understand the state of our victim networks and apply the steps described above.
-Using SGX-Step, we developed an attack application, which starts the enclave and interacts with it.
-We use this to record page accesses which we later interpret into activation function states.
-Using SGX-Step, we log all page accesses within the enclave (after a certain trigger page was accessed) and interpret the results as part of a post-processing step.
-Post-Processing
-Once Tensorflow Microlite ran within an enclave, we profiled the execution and observed call traces (via sgx-gdb) to understand execution flow.
-An example can be seen in LABEL:lst:sgx_expf_stack.
-This lets us build an intuition for how Tensorflow Microlite performed inference and which functions and pages are accessed.
+Results.
+Our attack fully recovered all the weights and biases for the neurons in the first layer of a Tensorflow Microlite network running inside an SGX enclave.
+After calibration, the full attack took 55 binary searches of the network per parameter to correctly recover each of the 1200 parameter in the first layer down to or beyond 3 decimal places of accuracy.
+This is in contrast to [7]’s budget of 100 queries per parameter.
+During the attack, each first-layer neurons’ parameters were recovered at around 99% accuracy (when compared to the ground truth values) by <36% of the way through the full search (around search depth 20).
+The average error is <1% at this point.
+The error continues to slowly improve until the search hits the limits of a 32-bit floating point number at around depth 55.
+These findings are further summarised in Appendix˜E, which shows how the achieved average and maximum error in the recovered weights and biases varies with the number of queries per neuron in Minsurance.
+The maximum error rates were higher in our full attack on Minsurance than we had seen in smaller scale testing.
+We speculate this is due to numerical precision issues when solving certain neurons, particularly ones with very small weights or perhaps a combination of very low and very high magnitude weights.
+During the binary search there may be an ideal magnitude for the non-dynamic input values based on the magnitude of the dynamic input.
+We consider this an area for future study and discuss a workaround below.
+As this network was substantially larger than those used in small scale testing, we developed two ideas to help accelerate the attack for the other case studies.
+Max Search Depth.
+We note that the attack can be sped-up by only binary searching to a certain depth of iterations.
+As can be seen in Appendix˜E, we already achieve an < 1% average error rate by depth 20.
+Based on the desired level of accuracy, an attacker can trade off between time and fidelity to the original network and choose to only search to some depth, rather than letting the binary search finish.
+We note that achieving sub 1% precision may not be as necessary as it seems; the widespread adoption of quantisation has shown that models can still be quite performant despite the precision loss incurred by the lossy conversion of hidden parameters to 8 bit integer representations.
+We also note that our sub 1% precision attack has complexity below that of [7], which also only targets classification networks and not continuous output networks like this one.
+Improved Method of Calibration.
+As described in Section˜IV-D1, the calibration phase saves us from wasting many search steps down from some large safe constant to the smallest values that will overflow/underflow the weights.
+In the previous equation, $S$ is therefore optimised by starting the search from the smallest possible point that still overflows/underflows our target.
+By automating this calibration process for each neuron, we ensure a search with the fewest steps possible.
+Though we first implemented calibration as a per-neuron process (as previously described) we realised when trying to attack this model that this step could be performed across an entire layer of neurons simultaneously.
+We can leverage the fact that our SGX-Step traces for a given input provide an instruction count measurement for all neurons simultaneously to perform a combined or input-centric calibration.
+We first create an input array for the network that is all zeros (one for each input).
+Next, we iterate through each index in the input, replacing the 0 with increasing powers of some constant, e.g., 10 and passing the array into the network.
+We then record when an input array has underflowed/overflowed all the neurons in the layer.
+This maximal value of the power of that constant, $D$, when multiplied by P1P-1 (the number of inputs) allows us to calculate $C$, the number of executions required to calibrate the whole network: $C=\,(P-1)\,\cdot D$.
+Note that $C$ is independent of NN and $S$.
+As $Q$ would likely be dominated by $N$ in a real network, this calibration phase should account for only a small fraction of the total queries in the full attack.
+By adding $C$ to $Q$ we get a good approximation for the number of executions required to recover the first layer and calibrate faster than via the original method.
+V-B Mmult: A Multi-Layer Model using sigmoid()
+Next, we show how to attack a more realistic activation function and recover partial information from deeper layers.
+Model.
+We now focus on a network trained to multiply two numbers with the following architecture: there are two inputs, followed by the first hidden layer which contains 4 sigmoid() neurons.
+Then there is a second hidden layer containing 8 sigmoid() neurons followed finally by the final layer with a single relu() neuron.
+Recovery.
+We follow the steps described in Section˜IV-D and Section˜IV-C to recover the first layer.
+The only difference is that we exploit a different CPU instruction count leakage; instead of targeting tlibc’s exponential function, we now attack the Tensorflow lite framework-level sigmoid() function as described in Section˜IV-B.
+This shows how our attack is adaptable to different and more realistic activation functions, as well as side channel leakages elsewhere in the ML inference pipeline.
+We then use the procedure described in Section˜IV-E to attack the second layer.
+Results.
+This recovery is summarised in Appendix˜F and is comparable to the results of our attack against Minsurance.
+We are also able to recover partial (signless) convergence sets for the neurons in the second layer.
+For example, we can recover 35 sets for the first neuron in the second layer.
+Given its 4 inputs and single bias term (5 total unknowns) these are many more convergence point sets than we would need for a solution.
+However, because we do not have sign information, we are unable to conclusively solve for the neurons’ parameters.
+V-C MMNIST: MNIST with sigmoid()
+Model.
+Lastly we discuss a more complex model which was trained on the MNIST dataset [54].
+Our MNIST case study is much more complex than the others before it featuring $28x28=784$ inputs and a first layer with $128$ neurons.
+Adding the biases, this network has $100,480$ hidden parameters in the first layer making it two orders of magnitude larger than Minsurance.
+For this study, we focus on only the first layer.
+Recovery.
+Here we combine parts of the recovery strategies of the first two case studies: a larger network which we attack using the framework-level sigmoid() vulnerability described in Section˜IV-B.
+As in the first model, we only attempt recovery of the first layer.
+Results.
+Because of the greater number of hidden parameters for this model, recovery using the methods described above would take millions of iterations and while feasible, not entirely practical.
+In Section˜V-A we discussed how the value $Q$ could be reduced by reducing the search depth, $S$.
+Another way to reduce $Q$ is to try to minimise the $P\,\cdot\,N\,=785\cdot 128=100480$ term.
+This is possible if we approach the binary search a different way, inspired by the input-centric calibration described previously.
+Rather than a neuron-centric regime (focusing on each neuron one at a time and ignoring the effect of the changing inputs on the other neurons in the same layer), we instead record the state of all the neurons in the layer simultaneously as they respond to changes to each input of the network in sequence.
+We note that in this model’s first layer, the weights are normally distributed meaning that there are clusters that can be easily searched together.
+This allows us to exploit the structure of the distribution of the hidden parameters, group weights with similar magnitudes, and combine searches allowing better than $P\,\cdot\,N\,$ performance.
+This also allows us to abandon searches in regions without any neurons.
+See the output of this process over an entire input in Appendix˜G and a visual in Appendix˜L.
+Note that this approach also integrates the input-centric calibration discussed above, which is completed for this input in 18 executions of the network.
+Also note that every iteration before depth 18 (where each neuron is effectively in its own search space) is a savings against the more naive approach described above.
+The output in Appendix˜G recurses to a minimum gap size of only $0.1$, so it does not have the same precision as the previous attack code example from Appendix˜C.
+Note that the output of the input-centric attack to depth $0.1$ (1050 executions) is roughly analogous in recovery precision to a depth $20$ neuron-centric attack (12820=2500128*20=2500 executions) meaning that we recover the same value using 42% of the executions.
+Extrapolating the full attack to maximum depth, we project that the neuron-centric approach would take $2,460,800$ executions (without calibration) and this input-centric approach would take $1,180,765$ executions with calibration, or 48%.
+The [7] attack requires a budget of 100 times the number of parameters in the network.
+If we only consider the first layer, this network would require a budget of (2828+1)128100=10,048,000(28*28+1)*128*100=10,048,000 queries (there are 28*28 weights between each input and each neuron, and we add one extra parameter for the bias term).
+Using the values above, we project a full depth scan of this layer in $1,180,765$ queries, which is 11% of the [7] budget.
+While we cannot report the accuracy of the full recovery, spot checks against ground truth values were very promising and could be made more (or less) precise by tuning the min gap parameter to stop the search when desired.
+VI Ecosystem Analysis
+We examined implementations of common activation functions in popular ML frameworks to assess if they exhibit input-dependent patterns.
+For this, we used a combination of static (Ghidra) and dynamic (gdb) analysis to explore the call stacks of these frameworks until an underlying maths function was reached.
+Our findings are summarised in Table˜II.
+There are several dimensions to our survey: first, we found that in all cases we tested, the low-level maths functions that underpin activation functions are not included in the ML frameworks themselves.
+Instead, they rely either on additional high performance maths libraries, e.g., sleef in the case of PyTorch, Eigen and DNNL in the case of Tensorflow and XNN in the case of Tensorflow Lite or call into standard libraries like tlibc (in the case of SGX) or glibc on a Linux system.
+In some cases, these frameworks do both for different functions, adding to the complexity of the analysis.
+The (in)security of an activation function can stem from the underlying library and implementation of these functions.
+For example, a call to sigmoid() that ends up in glibc’s highly input-dependent exp() in libm.so exposes the caller’s ML framework to vulnerability.
+Similarly, a call to the same function in a different framework that ends up in a highly optimised maths library like Eigen may be harder to exploit.
+However, as we note in Section˜IV-B and Section˜V-B, higher level framework code can also be the culprit of exploitable memory access patterns.
+The operational context may also play a roll in vulnerability; e.g., an embedded edge device may not have the hardware functionality for high performance or even floating point maths routines.
+The reason optimised maths libraries may be harder to exploit is their use of vectorised instructions (e.g., AVX on Intel or NEON on Advanced RISC Machines (ARM)) which discourage branching (since that can affect performance).
+For example, we observed that in PyTorch, softmax() was realised through calls into the sleef library, which at default optimisations compiled into AVX assembly code without memory access leakage.
+We note however that sleef incorporates a dispatch mechanism that calls into different implementations depending on the outcome of cpuid.
+Notably, if this library was used in an SGX enclave, where cpuid is typically implemented as an untrusted ocall, the adversary could spoof the cpuid result and force usage of the insecure non-AVX version.
+Curiously, for sigmoid(), PyTorch resorts to the standard library implementation of exp(), rendering the implementation insecure independent of cpuid.
+Tensorflow uses Eigen for all the functions we tested except for Exponential(), which calls directly into glibc.
+Finally it is worth restating that even though vectorised code is non-branching, it is not guaranteed to execute in constant time [27].
+Finally, Table˜II shows how the optimisation level has a direct impact on the security of relu() for Tensorflow Microlite.
+While the compiler default of -Os happens to be free of memory access leakage, this is more by accident than by design, which is not desirable in the case of libraries that handle sensitive data.
+Analogously, the security benefits of vectorised code against our attack is not a proactive security-related decision but a side effect of pursuing performance.
+TABLE II: Survey of vulnerability across various ML and standard libraries and for different optimisation levels.
+ML Framework RELU Exponential Sigmoid Softmax
+TFLiteMicro (SGX/tlibc, -O0)
+TFLiteMicro (SGX/tlibc, -Os)
+TFLiteMicro (glibc, -Os)
+TFLite (glibc)
+TensorFlow CPU (glibc)
+PyTorch (glibc) n.a. ✓*
+Legend: ✓: Secure ✗: Insecure due to data-dependent access pattern
+*: Only secure if cpuid indicates AVX support
+Vulnerability of relu().
+In addition to expf(), we also considered relu() and its implementation in Tensorflow Microlite, which relies on the std::max() function.
+When running inside of an enclave, the std::max() library call is provided by an implementation in SGXs tlibc library which can be seen in Appendix˜H.
+With the default settings (gcc 11.4.0 -Os with function marked inline), the above code was compiled into the instructions shown in Appendix˜I.
+When compiling with -O0, however, the function was assembled as shown in Appendix˜J.
+This is significant because when the compiler emits jmp-class instructions, it is possible (using sgx-gdb) to detect the single step difference and therefore discern between the cases where $input\leq 0$ and $input>0$.
+This is a threshold point between two instruction count classes about $0.0$, similar to the ones discussed previously in expf().
+Though the difference might only be a single CPU instruction (whether a jump was taken or not) this difference is detectable.
+Appendix˜I, on the hand, has no step count differences due to its use of the CMOVA instruction which combines a branch and a move into a single instruction.
+Whatever the outcome of the conditional, CMOVA always takes the same number of steps, removing the timing leak.
+Within our case study, relu() within Tensorflow Microlite was not vulnerable, but it was not secure by design either.
+This is worrisome because in contrast to vetted cryptographic code, ML libraries currently do not exhibit input-independent memory access and execution behaviour, and further the eventual security properties of high-level ML code may depend on compiler optimisations and other non-explicit behaviour as with relu() in some configurations.
+Without any specific safeguards taken, ML libraries are not intrinsically secured against the vulnerability discussed in this paper.
+We note that on x86, many floating point instructions can raise exceptions or otherwise exhibit operand-dependent timing as mentioned by [27] and (in a different context) [55], which might result in timing vulnerabilities even if branchless code is emitted [27, 55].
+We leave the exact measurement and exploitation of such timing side channels in real-world settings as an area for future research.
+Vulnerability of Framework Code.
+As discussed in Section˜IV-B, and demonstrated practically in Section˜IV-E, exploitable memory access pattern vulnerabilities can exist outside mathematical library code in ML frameworks themselves.
+This further supports our case that these frameworks are not suitable for enclave execution under a SCA advisory.
+Beyond the security properties of the libraries they use, security-conscious ML framework developers need to also consider their own code and if they are introducing exploitable timing side channels.
+This is especially true for more complex activation functions like softmax() which are not mathematical functions likely to be found in standard maths libraries.
+VII Discussion and Mitigations
+We show that input-dependent memory accesses and branching are prevalent in ML implementations, and that the necessity for secure programming practices is not widely understood in the ML community.
+Our PoCs shows that an attacker with a high-resolution controlled or side channel is able to exploit data-dependent memory accesses to recover model details, whereas prior software SCA typically only recovered hyperparameters such as model architecture [15].
+Though there are certain limitations, Minsurance and Mmult show precise weight recovery can be performed with at maximum 55 queries per weight, and at over 99% accuracy at less than half that number (plus calibration).
+Search depth can be parametrically adjusted to account for the desired accuracy.
+Ultimately, while out of scope for this paper, we believe that if an attacker can obtain precise-enough timing measurements, similar attacks may be possible outside of a TEE context, e.g., for virtualised cloud deployments.
+Cache attacks with high temporal resolution, e.g., Prime+Scope [56], could be enough to recover traces similar to SGX-Step, though attacks would likely require modifications to deal with increased noise.
+Regardless of whether ML is being performed in enclaves, we argue that hence, ML library developers should pay attention to the full implementation attack surface including library and framework code.
+Even if e.g., relu() implementations were often secure in the surveyed libraries, this largely relied on certain compiler optimisations and was not ‘by design’.
+Similar to cryptographic libraries, ML libraries should consider hardening their codebase against software side channels such as data-dependent memory access patterns.
+Security and privacy-sensitive algorithms, to which ML now belongs, must be robust by design, regardless of context, platform, or (mis)use of library interfaces by a programmer.
+We briefly explored quantised networks as an extension to our work.
+With fewer bits representing each hidden parameter, we reasoned binary searches might encounter discrete boundaries in fewer iterations.
+We converted our MNIST model using Tensorflow’s full integer quantisation to a Tensorflow lite model with 8bit integer representations of its weights and biases and captured traces using our SGX-Step setup (an inset of which is visualised in Appendix˜M).
+Aside from the additional (de)quantisation phases during the network’s execution (not pictured), the patterns of page accesses and timings were very different to previous traces, owning to the different functions / processes used to perform inference on fixed point values.
+Though we did not further investigate the access patterns for this paper, we noticed evidence of non-constant time behaviour in the evaluation of neighbouring neurons suggesting non-constant time effects.
+Countermeasures.
+As with the work done in cryptography libraries, there are established steps to avoid data-dependent branching and memory accesses, e.g., by relying on bitwise logical operations, constant-time conditional moves, and other related techniques.
+For x86, as discussed for the sleef library, using vectorised SIMD code can also mitigate the issue, though needs to be thoroughly vetted for other side channels like floating point timing [55, 27].
+At present, akin to a “chosen plaintext” attack for cryptographic algorithms, our attacks entails sending arbitrarily large or small floating point numbers into the network.
+Any input normalisation might mitigate the fast calibration and binary search approaches described in this paper, only enabling partial recovery through the grid search technique.
+Similarly, as the attack requires a large number of queries to the network, rate-limiting or detection of “suspicious” usage would render exploitation harder in practice.
+However, again, ML security should in our view not rest on developers having to normalise inputs or limit performance, so these measures are supplementary to a securely developed library.
+Finally, researchers have explored defenses at the level of the TEE itself to frustrate (but not fully eliminate) side-channel leakage exploitation.
+Initial approaches [57, 58, 59] focused on detecting suspicious interrupt rates as a side-effect of an ongoing controlled-channel attack, which may be error-prone and suffer from false positives.
+Alternatively, custom oblivious RAM solutions [60, 61, 62] may probabilistically hide secret-dependent enclave memory accesses, however this is at the cost of prohibitive performance overheads.
+Most prominently, AEX-Notify [41] is a recent Intel SGX hardware-software extension that aims to eliminate SGX-Step’s single-stepping capabilities by prefetching selected application pages.
+Notably, AEX-Notify is an opt-in feature only available in recent CPUs.
+Enclaves that did opt-in would preclude the ability to gather precise instruction counts through single-stepping.
+However, AEX-Notify explicitly does not prevent general information leakage through page faults and, hence, cannot fully mitigate our attacks.
+While specialised, compiler-assisted defenses [34, 63] have been proposed to mitigate page-fault leakage on off-the-shelf SGX platforms, these approaches incur prohibitive performance overheads, and their applicability to general-purpose ML workloads remains unexplored.
+Limitations.
+We require the victim to run in a co-located TEE in order to use our controlled channel, which is a different requirement than previous model stealing attacks, that require chosen inputs and the output distribution of the inference pass.
+Despite the recovery of the first layer of neurons fundamentally undermining SGX’s promise to secure the entire computation, our attack is limited in recovering parameters from deeper layers, depending on the model architecture.
+This is mostly due to the (limited) expressive power of some common activation functions as noted in Tables˜I and IV-E.
+The grid search is less efficient than the binary search used for the first layers, and depending on the magnitude of the weights, small search steps might be needed to discover thresholds.
+An adaptive approach could deepen the sweep region until a threshold is hit; however, small weights likely contribute little to the network and may be replaced by a static value to produce an equivalent (non-identical) network.
+Finally, we focused on input-dependent memory access patterns and instruction counts in this paper, but did not take input-dependent latency of single instructions into account.
+For example, floating point multiplication is not necessarily guaranteed to execute in constant time [27].
+This could render vectorised code like the implementation from sleef vulnerable to a modified version of our attack: while different inputs might have identical instruction counts (as reported by SGX-Step), the actual runtime might be different and be observable with other side channels like interrupt latency [37].
+VIII Conclusion
+The recent surge in ML applications necessitates an in-depth understanding of new attack vectors and defenses.
+In this respect, confidential computing architectures, and in particular Intel SGX, have gained significant traction in recent years and hold the potential to securely outsource critical ML computations to pervasive untrusted remote cloud platforms.
+However, we show that SGX-protected ML implementations exhibit subtle side-channel vulnerabilities that allow the extraction of precise model parameters in our PoC attacks.
+In the wider perspective, our work may generalise to other TEEs and highlights the issue of input-dependent memory accesses that are prevalent in today’s ML implementations, suggesting that security-critical ML applications should consider adopting coding practices studied in the cryptographic community.
+Acknowledgments
+The PhD of Jesse Spielman is partially funded by a donation from Intel.
+This research was further partially funded by the Engineering and Physical Sciences Research Council (EPSRC) under grants EP/X03738X/1 and EP/R012598/1.
+This research is partially funded by the Research Fund KU Leuven, and by the Cybersecurity Research Program Flanders.
+Appendix A Implementation of sigmoid() (called logistic()) in Tensorflow lite
+1inline void Logistic(const RuntimeShape& input_shape, const float* input_data,
+2 const RuntimeShape& output_shape, float* output_data) {
+3 const float cutoff_upper = 16.619047164916992188f;
+4 const float cutoff_lower = -9.f;
+5
+6 const int flat_size = MatchingFlatSize(input_shape, output_shape);
+7
+8 // [comments removed]
+9
+10 for (int i = 0; i < flat_size; i++) {
+11 float val = input_data[i];
+12 float result;
+13 if (val > cutoff_upper) {
+14 result = 1.0f;
+15 } else if (val < cutoff_lower) {
+16 result = std::exp(val);
+17 } else {
+18 result = 1.f / (1.f + std::exp(-val));
+19 }
+20 output_data[i] = result;
+21 }
+22}
+Listing 2: Implementation of Logistic() in Tensorflow lite
+Appendix B expf() call stack for Tensorflow Microlite neuron
 10x7ffff762b460 in expf()
@@
 140x555555557a24 in main()
-Listing 2: Simplified call stack for expf()-based neuron in Tensorflow Microlite.
-Unlike sgx-gdb, SGX-Step only offers a relatively coarse-grained spatial resolution of 4 KiB memory pages, which prevents production enclave attackers from directly identifying the specific symbol (e.g., function or data) accessed by the enclave code.
-However, SGX-Step’s precise instruction-level temporal resolution enables the annotation of coarse-grained page-access traces with the exact number of instructions executed on each page.
-As illustrated in Figure 4, this capability allows to accurately reconstruct the inference execution structure, even without knowledge of the lower 12-bit page offset of the secret enclave instruction pointer (ERIP).
-By comparing the dotted red line which depicts the true ERIP values available to sgx-gdb and the black line which shows the reduced page-level granularity, one can see how little is lost.
-It is worth nothing that each page can contain multiple functions and so it is difficult to assign a precise “meaning” to each page without context.
-The instructions executed on one page could pertain to two very different activities at different points in execution.
-However, the combination of coarse-grained page-access patterns along with the exact amount of instructions executed per page, annotated in black text, clearly suffices to identify individual function calls or data accesses.
-In practice, this means that even if multiple symbols are co-located on the same 4 KiB page, we can still interpret the logs.
-We also know precisely the length (or length range) of different runs at different points during the execution, i.e., we can also tell if a trace is invalid and immediately reject it.
-Figure 4: Visualization of SGX-Step’s instruction-granular page access trace.
-Inset of one layer’s 16 neurons.
-Note that the intra-page instruction pointer (RIP) values highlighted in dotted red are only displayed for reference visualization purposes here and are not used by the final attack.
-By studying inference traces from the perspective of both sgx-gdb and SGX-Step (raw traces or visualizations like those shown in Figure 4), clear patterns emerge.
-For example, the function ExpEval() is located on a different page to expf() (each 0x1000 bytes is a new page) and we can plainly interpret the movement between them as the higher level neuron code invoking the lower level expf() math function 16 times (once for each neuron in this layer).
-The 18 instruction count long runs at the extreme left and right on page 0x2c000 seem to correspond to the start and end of a layer, which makes parsing each layer trivial.
-We note that the step counts provided by sgx-gdb and SGX-Step may vary slightly (but deterministically) due to the documented effects of macro-op fusion [moghimi2020copycat].
-The final part of our tooling is a Python script and library which further abstracts and manages invocations of the enclave program.
-Because each neuron is accessed in order and each return occurs in sequence, it should be possible to monitor each neuron’s step count independently.
-In order to orchestrate the full attack, we use this script to interpret the log output by SGX-Step into a per-neuron state.
-Even though we receive the state for every neuron in each log, we only focus on one at a time.
-An example output of the running attack is shown in LABEL:lst:recovery_attack_output in Appendix A.
-6.2 Multi-Layer Model with sigmoid()
-Model
-We now focus on a “multiplication” network (trained to multiply numbers), a small network with the following architecture: there are two inputs, followed by the first hidden layer which contains 4 sigmoid neurons.
-Then there is a second hidden layer containing 8 sigmoid neurons followed finally by the final layer with a single relu neuron.
-Recovery
-We follow the steps described in Section 5.2 and Section 6.1 to recover the first layer using the exact same attack code.
-The only difference is that we exploit a different CPU instruction count leakage; instead of targeting tlibc’s exponential function, we now attack the Tensorflow lite framework-level sigmoid function as described in Section 4.1.
-This shows how our attack is adaptable to different and more realistic activation functions, as well as side channel leakages elsewhere in the ML inference pipeline, cf.
-Section 7.1.1.
-7 Results
-We integrated our victim enclave, performing inference using our Tensorflow Microlite network, with SGX-Step for side-channel observations and our attack script for weight and bias recovery.
-In Section 7.1 we discuss the results from the attack against the proof of concept regression network.
-In Section 7.2 we discuss how our findings relate to other activation functions in other ML Frameworks.
-As a part of that, we more explicitly consider relu() (Section 7.2.1) and framework code itself (Section 7.2.2).
-7.1 End-to-End Attack Case Study for Tensorflow Microlite Enclave
-Our attack was able to fully recover all the weights and biases for the neurons in the first layer of a Tensorflow Microlite network running inside an SGX enclave.
-After calibration, the full attack took 55 binary searches of the network per parameter to recover each of the 1200 parameter in the first layer down to or beyond 3 decimal places of accuracy.
-During the attack, each first-layer neurons’ parameters were recovered at around 99% accuracy (when compared to the ground truth values) by <36% of the way through the full search (around search depth 20).
-The average error is <1% at this point.
-The error continues to slowly improve until the search hits the limits of a 32-bit floating point number at around depth 55.
-These findings are further summarised in Table II, which shows how the achieved average and maximum error in the recovered weights and biases varies with the number of queries per neuron in our 11 input regression ‘victim’.
-The maximum error rates were higher in our full attack on the regression network than we had seen in smaller scale testing.
-We speculate this is due to numerical precision issues when solving certain neurons, particularly ones with very small weights or perhaps a combination of very low and very high magnitude weights.
-During the binary search there may be an ideal magnitude for the non-dynamic input values based on the magnitude of the dynamic input.
-We consider this an area for future study and discuss a workaround below.
-The speed (e.g., number of invocations) is dependent on the architecture of the network (the number of parameters to recover) but also some factors the attacker can control such as the depth of the binary search.
-For example, the attack can be sped-up by only binary searching to a certain depth of iterations.
-As can be seen in Table II, we already achieve an < 1% average error rate by depth 20.
-Based on the desired level of accuracy, an attacker can choose to only search to some depth, rather than letting the binary search finish.
-We note that achieving sub 1% precision may not be as necessary as it seems.
-As the widespread adoption of quantization has shown, some models are still accurate with the precision loss due to the conversion to 8 bit integers.
-In general, the total number $Q$ of queries to recover the first layer can be computed as:
-$Q=\,P\,\cdot\,N\,\cdot S$
-where $P$ is the number of hidden parameters (in the first layer, e.g., one weight per input plus one for the bias) to solve for, $N$ is the number of neurons in the first layer, and $S$ is the maximum number of steps in the binary search.
-The $P$ value correlates to the minimum number of equations we require to solve a system of equations: one for each unknown.
-We can compensate for noise or other errors by adding extra equations to produce better solutions at the cost of more queries.
-In testing, finding 3 additional convergence sets helped reduce the maximum error by 28% at depth 25 compared to an attack where precisely $P$ convergence sets were generated (this is visualised in Figure 5 in Appendix C).
-It might be possible with future study to detect particularly weak systems of equations analytically and automatically add extra convergence sets to improve the results dynamically.
-Secondly, we note the effect Calibration (Section 5.2.1) has on the performance of the attack.
-We can ensure we start the binary search in an overflow/underflow state by choosing a very high magnitude starting value for our dynamic input at the cost of having to waste many queries binary searching down to the range most hidden parameters occupy near 0.
-The performance of $S$ is therefore optimised by starting the search from the smallest possible point that still overflows our target.
-Rather than starting from an extreme value like MAX_FLOAT, a better alternative is to apply the heuristic that a weight likely has a magnitude between -1 and 1 and see if 100 is enough to overflow it.
-If not, we try 1000, then 10,000 and so on.
-By automating this calibration process for each neuron, we ensure a search with the fewest steps possible.
-In our experiments we implemented calibration as a per-neuron process but we realised that this step could be performed across an entire layer simultaneously.
-In contrast to search phases of the attack, we can leverage the fact that we receive an instruction count measurement for all neurons simultaneously for each execution of the network.
-We create an input to the network that is all 0s and then iterate through each index in the input, replacing that 0 with increasing powers of 10 until we’ve under/overflowed the smallest weighted neuron in the layer.
-This maximal power of 10, $D$, when multiplied by $P-1$ (the number of inputs) allows us to calculate $C$, the number of executions required to calibrate the whole network.
-$C=\,(P-1)\,\cdot D$
-We note that $C$ is totally orthogonal to $N$ and $S$.
-Because $Q$ would likely be dominated by $N$ in a real network, this calibration phase should account for only a small fraction of the total queries in the full attack.
-By adding $C$ to $Q$ we get a good approximation for the number of executions required to recover the first layer.
-TABLE II: Average and Maximum Error for different search iterations of 11 input Regression Network.
-Iterations Average error Max. error
-5 6.029863100281893 844.0875066255469
-10 2.8364600666971387 786.2672913954896
-15 0.27165280197604547 156.03829543970028
-20 0.008040720460419709 4.501496529005743
-25 0.0005365116007451519 0.25501229735941244
-30 0.0003891028780509273 0.1506828257822338
-35 0.00038239633382359514 0.14650184342297062
-40 0.000382652272853835 0.14664602195446264
-45 0.00038265156035156687 0.14664546962120745
-50 0.000382651519816129 0.14664550955005118
-55 0.00042219596795115005 0.14664550840498652
-[…]
-7.1.1 Results for Multiplication network
-This recovery is summarised in Table III and is comparable to the results of our attack against the other model.
-We are also able to recover partial (signless) convergence sets for the neurons in the second layer.
-For example, we can recover 35 sets for the first neuron in the second layer.
-Given its 4 inputs and single bias term (5 total unknowns) these are many more convergence point sets than we would need for a solution.
-However, because we do not have sign information, we are unable to conclusively solve for the neurons’ parameters.
-TABLE III: Average and Maximum Error for different search iterations of 2-input Multiplication Network.
-Iterations Average error Max. error
-5 5.104921006404077 17.916865999472293
-10 1.6756457515509167 17.51883507815715
-15 1.5295341684026322 17.520958313634686
-20 0.06962602190494313 0.7988771642785077
-25 0.002053382484092837 0.023557097950552608
-30 4.133456316032444e-05 0.0004519024401035132
-35 3.313471981702896e-05 0.00036192843021964904
-40 3.2159960239434036e-05 0.0003506233064154429
-45 3.215762491123297e-05 0.00035059400594272816
-50 3.215754896549253e-05 0.0003505930825000725
-55 3.21575691725288e-05 0.0003505933114027471
-7.2 Activation Function Safety Ecosystem Study
-In order to extend our findings and consider the breadth of the issue, we examined implementations of common activation functions to determine in popular ML frameworks to assess if they exhibit input-dependent memory access patterns.
-For this, we used a combination of static (Ghidra) and dynamic (gdb) analysis to explore the often complex call stacks of these frameworks until an underlying math function was reached.
-Our findings are summarised in Table IV.
-There are several dimensions to our survey: first, we found that in all cases we tested, the low-level math functions that underpin activation functions are not included in the ML frameworks themselves.
-Instead, they rely either on additional high performance math libraries, e.g., sleef in the case of PyTorch, Eigen and DNNL in the case of Tensorflow and XNN in the case of Tensorflow Lite or call into standard libraries like tlibc (in the case of SGX) or glibc on a Linux system.
-In some cases, these frameworks do both for different functions, adding to the complexity of the analysis.
-The (in)security of an activation function can stem from the underlying library and implementation of these functions.
-For example, a call to sigmoid() that ends up in glibc’s highly input-dependent exp() in libm.so exposes the caller’s ML framework to vulnerability.
-Similarly, a call to the same function in a different framework that ends up in a highly optimised math library like Eigen may be harder to exploit.
-However, as we note in Section 4.1 and Section 6.2, higher level framework code can also be the culprit of exploitable memory access patterns.
-The hardware context may also play a roll in vulnerability; an Internet of Things (IoT) device performing inference on an embedded system for example may not have the hardware functionality or enough space for advanced math routines beyond a basic standard library.
-The reason optimised math libraries may be harder to attack is that their use of vectorised instructions (for example AVX on Intel or NEON on Advanced RISC Machines (ARM)) which discourage branching (since that can affect performance).
-For example, we observed that in PyTorch, softmax() was realised through calls into the sleef library, which at default optimisations compiled into AVX assembly code without memory access leakage.
-We note however that sleef incorporates a dispatch mechanism that calls into different implementations depending on the outcome of cpuid.
-Notably, if this library was used in an SGX enclave, where cpuid is typically implemented as an untrusted ocall [^3] https://github.com/intel/linux-sgx/blob/80a662/common/inc/sgx_tstdc.edl, the adversary could spoof the cpuid result and force usage of the insecure non-AVX version.
-Curiously, for sigmoid(), PyTorch resorts to the standard library implementation of exp(), rendering the implementation insecure independent of cpuid.
-Tensorflow uses Eigen for all the functions we tested except for Exponential(), which calls directly into glibc.
-Finally it is worth restating that even though vectorised code is non-branching, it still may not be constant time, which we discuss further in Section 8.3.
-Finally, Table IV shows how the optimisation level has a direct impact on the security of relu() in the case of Tensorflow Microlite as further discussed in Section 7.2.1.
-While the compiler default of -Os happens to be free of memory access leakage, this is more by accident than by design, which is not desirable in the case of libraries that handle sensitive data.
-Analogously, the security benefits (with regard to the attack we have demonstrated) of vectorised code is not a proactive security-related decision but a side effect of pursuing performance.
-TABLE IV: Survey of vulnerability across various ML and standard libraries and for different optimisation levels.
-ML Framework RELU Exponential Sigmoid Softmax
-TFLiteMicro (SGX/tlibc, -O0)
-TFLiteMicro (SGX/tlibc, -Os)
-TFLiteMicro (glibc, -Os)
-TFLite (glibc)
-TensorFlow CPU (glibc)
-PyTorch (glibc) n.a. ✓*
-Legend: ✓: Secure ✗: Insecure due to data-dependent access pattern
-*: Only secure if cpuid indicates AVX support
-7.2.1 Vulnerability of relu()
-In addition to expf(), we also considered relu() and its implementation in Tensorflow Microlite, which relies on the std::max() function.
-When running inside of an enclave, the std::max() library call is provided by an implementation in SGXs tlibc library which can be seen in LABEL:lst:max_in_c).
-1template <class _T1>
-2struct __less<_T1, _T1>
-3{
-4 _LIBCPP_INLINE_VISIBILITY _LIBCPP_CONSTEXPR_AFTER_CXX11
-5 bool operator()(const _T1& __x, const _T1& __y) const {return __x < __y;}
-6};
-7
-8template <class _Tp, class _Compare>
-9inline _LIBCPP_INLINE_VISIBILITY _LIBCPP_CONSTEXPR_AFTER_CXX11
-10const _Tp&
-11max(const _Tp& __a, const _Tp& __b, _Compare __comp)
-12{
-13 return __comp(__a, __b) ? __b : __a;
-14}
-Listing 3: std::max() source code from sgxsdk/include/libcxx/algorithm.
-With the default settings (-Os passed to the compiler, gcc 11.4.0, and the function marked as inline), the above code was compiled into branchless (e.g., UCOMISS) instructions as shown in LABEL:lst:max_in_asm.
-In contrast, when compiling with -O0, the function was assembled with conditional branches as shown in LABEL:lst:max_in_asm_o0.
-This is significant because when the compiler emitted jmp-class instructions, it was possible (using sgx-gdb) to detect the single step difference and therefore discern between the cases where $input\leq 0$ and $input>0$.
-This is a threshold point between two instruction count classes about $0.0$, similar to the ones discussed previously in expf().
-Though the difference might only be a single CPU instruction (whether a jump was taken or not) this difference is detectable.
-100100a04 MOVSS XMM0,dword ptr [param_2]
-200100a08 UCOMISS XMM0,dword ptr [param_1]
-300100a0b MOV RAX,param_1
-400100a0e CMOVA RAX,param_2
-500100a12 RET
-Listing 4: Resulting assembly for std::max() at -Os with gcc 11.4.0.
-1001019c9 PUSH RBP
-2001019ca MOV RBP,RSP
-3001019cd MOV qword ptr [RBP + local_10],param_1
-4001019d1 MOV qword ptr [RBP + local_18],param_2
-5001019d5 MOV RAX,qword ptr [RBP + local_10]
-6001019d9 MOVSS XMM1,dword ptr [RAX]
-7001019dd MOV RAX,qword ptr [RBP + local_18]
-8001019e1 MOVSS XMM0,dword ptr [RAX]
-9001019e5 COMISS XMM0,XMM1
-10001019e8 JBE 001019f0
-11001019ea MOV RAX,qword ptr [RBP + local_18]
-12001019ee JMP 001019f4
-13001019f0 MOV RAX,qword ptr [RBP + local_10]
-14001019f4 POP RBP
-15001019f5 RET
-Listing 5: Resulting assembly for std::max() at -O0 with gcc 11.4.0.
-Within the our case study, relu() within Tensorflow Microlite was not vulnerable, but it was not secure by design either.
-This is worrisome because in contrast to vetted cryptographic code, ML libraries currently do not exhibit input-independent memory access and execution behavior, and further the eventual security properties of high-level ML code may depend on compiler optimizations and other non-explicit behavior as with relu() in some configurations.
-Without any specific safeguards taken, ML libraries are not intrinsically secured against the vulnerability discussed in this paper.
-As a sidenote, though not considered in this paper, we note that on x86, many floating point instructions can raise exceptions or otherwise exhibit operand-dependent timing as mentioned by [Gongye20] and (in a different context) [Kohlbrenner17], which might result in timing vulnerabilities even if branchless code is emitted.
-We leave the exact measurement and exploitation of such timing side channels in real-world settings as an area for future research.
-7.2.2 Vulnerability of Framework Code
-As discussed in Section 4.1, and demonstrated practically in Section 5.3, exploitable memory access pattern vulnerabilities can exist outside mathematical library code in ML frameworks themselves.
-This further supports our case that these frameworks are not suitable for enclave execution under a Side-Channel Analysis (SCA) advisory.
-Beyond the security properties of the libraries they use, security-conscious ML framework developers need to also consider their own code and if they are introducing exploitable timing side channels.
-This is especially true for more complex activation functions like softmax() which are not common mathematical functions likely to be found in standard math libraries.
-8 Discussion and Mitigations
-First, we demonstrated that SGX-protected ML implementations can exhibit vulnerabilities due to oversights in their implementation, leading to extraction of precise model weights and biases.
-Our proof-of-concept shows that an attacker with a high-resolution controlled or side channel is able to exploit data-dependent memory accesses to recover model details, whereas prior software side channel attacks typically only recovered hyperparameters such as model architecture [yan_cache_2020].
-Though there are certain limitations, precise weight recovery can be performed with at maximum 55 queries per weight, and at over 99% accuracy at less than half that number (plus calibration).
-Speed can be parametrically adjusted to account for the desired accuracy.
-While our attack demonstration focused on the popular Intel SGX architecture, the underlying insights and weight-recovery techniques are applicable to other TEEs.
-Notably, the required side-channel primitives offered by SGX-Step, i.e., precise single-stepping and page accesses, have been demonstrated on alternative VM-based TEEs like AMD SEV [Wilke_Wichelmann_Rabich_Eisenbarth_2023] and Intel TDX [Wilke24, shome2024tdexit-notify].
-Ultimately, while not in the scope of this paper, we believe that if an attacker can obtain precise-enough timing measurements, similar attacks may be possible outside of a TEE context, e.g., for virtualised cloud deployments.
-Cache attacks with high temporal resolution, e.g., Prime+Scope [Purnal21], could be applicable to recover the access patterns similar to SGX-Step, though attacks would likely require modifications to deal with increased noise.
-We further show that input-dependent memory accesses and branching are prevalent in ML implementations, and that the necessity for secure programming practices is not widely understood in the ML community.
-Even if the attack described in this paper currently does not apply to non-enclaved targets, we argue that ML library developers should pay attention to the implementation attack surface.
-For example, even if relu() implementations were often secure in the surveyed libraries, this largely relied on certain compiler optimizations and was not ‘by design’.
-Similar to cryptographic libraries, ML libraries should hence consider hardening their codebase against software side channels such as data-dependent memory access patterns.
-Security and privacy-sensitive algorithms, to which ML now belongs, must be robust by design, regardless of context, platform, or (mis)use of library interfaces by a programmer.
-8.1 Countermeasures
-As with the work done in cryptography libraries, there are established steps to avoid data-dependent branching and memory accesses, for example by relying on bitwise logical operations, constant-time conditional moves, and other related techniques.
-For x86, as discussed for the sleef library, using vectorised SIMD code can also mitigate the issue, though needs to be thoroughly vetted for other side channels like floating point timing [Kohlbrenner17, Gongye20].
-At present, akin to a “chosen plaintext” attack for cryptographic algorithms, our attacks entails sending arbitrarily large or small floating point numbers into the network.
-Any input normalization or quantisation might mitigate the fast calibration and binary search approaches described in this paper, only enabling partial recovery through the grid search technique.
-Similarly, as the attack requires a large number of queries to the network, rate-limiting or detection of “suspicious” usage would render exploitation harder in practice.
-However, again, ML security should in our view not rest on developers having to normalise inputs or limit performance, so these measures are supplementary to a securely developed library.
-Finally, researchers have explored defenses at the level of the TEE itself to not fully eliminate side-channel leakage, but frustrate exploitation in practice.
-Initial approaches [shih2017tsgx, gridin2020dejavu, oleksenko2018varys] focused on detecting suspicious interrupt rates as a side-effect of an ongoing controlled-channel attack, which may be error-prone and suffer from false positives.
-Alternatively, custom oblivious RAM solutions [aga2019invisipage, zhang2020klotski, brasser2019dr] may probabilistically hide secret-dependent enclave memory accesses, however this is at the cost of prohibitive performance overheads.
-Most prominently, AEX-Notify [constable2023aexnotify] is a recent Intel SGX hardware-software extension that aims to eliminate SGX-Step’s single-stepping capabilities by prefetching selected application pages.
-Notably, AEX-Notify is an opt-in feature that is only available in recent CPUs.
-With regards to our attacks, victim enclaves that opt-in to AEX-Notify would preclude the ability to gather precise instruction counts through single-stepping.
-However, AEX-Notify explicitly does not prevent general information leakage through page faults and, hence, cannot fully mitigate our attacks.
-8.2 Limitations
-There are a certain limitations to our proof of concept.
-We require the victim to run in a co-located TEE in order to use our controlled channel, which is a different requirement than previous model stealing attacks, which require chosen inputs and the output distribution of the inference pass.
-Despite the recovery of the first layer of neurons fundamentally undermining SGX’s promise to secure the entire computation, our attack is limited in recovering parameters from deeper layers, depending on the model architecture.
-This is mostly due to the (limited) expressive power of some common activation functions as noted in Table I and detailed in Section 5.3.
-Besides, the used grid search is less efficient than binary search for the first layers, and depending on the magnitude of the weights, small search steps might be needed to discover thresholds.
-An adaptive approach could be used where the sweep region is deepened until a threshold is hit.
-We note however that small weights likely contribute little to the network, and thus could be replaced by a static value during the recovery process, producing an equivalent (though not identical) network.
-8.3 Future Work
-To efficiently recover missing parameters from deeper layers, if the output of the network is available, our attack could be combined with a Tramèr-style attack.
-After a partial parameter recovery, the obtained parameters would be placed into an untrained network with identical architecture, and be locked during training as in transfer learning.
-Then, the remaining neurons can be trained using examples generated from input/output pairs from the source network.
-In particular, image processing and similar networks contain a large proportion of their parameters in the early layers (and have an overall funnel shape), hence, this might result in a more precise recovery compared to a brute-force approach.
-Besides, we demonstrated our attack approach for simple FNNs.
-It would be interesting to expanding the attack to more complex architectures like Convolutional Neural Networks or transformers, using a library OS like Gramine [^4] https://github.com/gramineproject/gramine to run the relevant code in an SGX enclave.
-To improve the performance of our attack, a weighted binary search could reduce the number of search steps, as could caching some shared state such as the during the calibration step.
-Finally, we focused on input-dependent memory access patterns and instruction counts in this paper, but did not take input-dependent latency of single instructions into account.
-For example, the UCOMISS instruction discussed in Section 7.2.1 is not necessarily guaranteed to execute in constant time [Gongye20], and can raise floating point exceptions.
-This could render vectorised code like the implementation from sleef vulnerable to a modified version of our attack: while different inputs might have identical instruction counts (as reported by SGX-Step), the actual runtime might be different and be observable with other side channels like IRQ latency [VanBulck18].
-9 Conclusion
-The recent surge in ML applications necessitates an in-depth understanding of new types of attack vectors and defenses.
-In this respect, confidential computing architectures, and in particular Intel SGX, have gained significant traction in recent years and hold the potential to securely outsourcing critical ML computations to pervasive untrusted remote cloud platforms.
-This paper showed, however, that SGX-protected ML implementations exhibit subtle side-channel vulnerabilities that allow to extract precise model weight and biases in practical proof-of-concept attacks.
-In the wider perspective, our work may generalise to other TEEs and highlights the issue of input-dependent memory accesses that are prevalent in today’s ML implementations, suggesting that security-critical ML applications should consider adopting coding practices studied in the cryptographic community.
-Acknowledgments
-The PhD of Jesse Spielman is partially funded by a donation from Intel.
-This research was further partially funded by the Engineering and Physical Sciences Research Council (EPSRC) under grants EP/X03738X/1 and EP/R012598/1.
-This research is partially funded by the Research Fund KU Leuven, and by the Cybersecurity Research Program Flanders.
-[…]
-Appendix A Example Attack Output
+Listing 3: Simplified call stack for expf()-based neuron in
+Tensorflow Microlite.
+Appendix C Neuron-Centric Attack Output
 1 [ Neuron 2 ]
@@
 41 Checkpoint saved!
-Listing 6: Example output from our attack tool finding convergence points.
-Appendix B Implementation of sigmoid() (called logistic()) in Tensorflow lite
-1inline void Logistic(const RuntimeShape& input_shape, const float* input_data,
-2 const RuntimeShape& output_shape, float* output_data) {
-3 const float cutoff_upper = 16.619047164916992188f;
-4 const float cutoff_lower = -9.f;
-5
-6 const int flat_size = MatchingFlatSize(input_shape, output_shape);
-7
-8 // [comments removed]
-9
-10 for (int i = 0; i < flat_size; i++) {
-11 float val = input_data[i];
-12 float result;
-13 if (val > cutoff_upper) {
-14 result = 1.0f;
-15 } else if (val < cutoff_lower) {
-16 result = std::exp(val);
-17 } else {
-18 result = 1.f / (1.f + std::exp(-val));
+Listing 4: Example output from our neuron-centric attack tool
+finding convergence points.
+Appendix D Layer Unwrapping
+$act(input*W+b)$ $=target$
+$input*W+b$ $=act^{-1}(target)$
+$input*W$ $=act^{-1}(target)-b$
+$input*\cancel{W*W^{-1}}$ $=(act^{-1}(target)-b)*W^{-1}$
+$input$ $=(act^{-1}(target)-b)*W^{-1}$
+Figure 5: Note that it is also possible to do this without matrix inversion for greater numerical efficiency, e.g., using numpy.linalg.lstsq(W, act_inv(target)-b).
+Appendix E Minsurance Results
+TABLE III: Average and Maximum Error for different search
+iterations of Minsurance
+Iterations Average error Max. error
+5 $6.029863100281893$ $844.0875066255469$
+10 $2.836460066697138$ $786.2672913954896$
+15 $0.271652801976045$ $156.0382954397002$
+20 $0.008040720460419$ $4.5014965290057$
+25 $0.000536511600745$ $0.2550122973594$
+30 $0.000389102878050$ $0.1506828257822$
+35 $0.000382396333823$ $0.1465018434229$
+40 $0.000382652272853$ $0.1466460219544$
+45 $0.000382651560351$ $0.1466454696212$
+50 $0.000382651519816$ $0.1466455095500$
+55 $0.000422195967951$ $0.1466455084049$
+Appendix F Mmult Results
+TABLE IV: Average and Maximum Error for different search
+iterations of Mmult
+Iterations Average error Max. error
+5 $5.104921006404077$ $17.91686599947229$
+10 $1.675645751550916$ $17.51883507815715$
+15 $1.529534168402632$ $17.52095831363468$
+20 $0.069626021904943$ $0.79887716427850$
+25 $0.002053382484092$ $0.02355709795055$
+30 $0.000041334563160$ $0.00045190244010$
+35 $0.000033134719817$ $0.00036192843021$
+40 $0.000032159960239$ $0.00035062330641$
+45 $0.000032157624911$ $0.00035059400594$
+50 $0.000032157548965$ $0.00035059308250$
+55 $0.000032157569172$ $0.00035059331140$
+Appendix G Input-Centric Attack Output
+1Equation 3 / 785 / Other Value: 0.017745 / Input idx: 2
+2Calibrating...
+3Calibrated in: 18 executions to 131072
+4iteration: 0 / gaps: 1 / spans: 131072.0
+5iteration: 1 / gaps: 2 / spans: 65536.0
+6iteration: 2 / gaps: 2 / spans: 32768.0
+7iteration: 3 / gaps: 2 / spans: 16384.0
+8iteration: 4 / gaps: 3 / spans: 8192.0
+9iteration: 5 / gaps: 5 / spans: 4096.0
+10iteration: 6 / gaps: 7 / spans: 2048.0
+11iteration: 7 / gaps: 9 / spans: 1024.0
+12iteration: 8 / gaps: 12 / spans: 512.0
+13iteration: 9 / gaps: 17 / spans: 256.0
+14iteration: 10 / gaps: 25 / spans: 128.0
+15iteration: 11 / gaps: 34 / spans: 64.0
+16iteration: 12 / gaps: 43 / spans: 32.0
+17iteration: 13 / gaps: 55 / spans: 16.0
+18iteration: 14 / gaps: 74 / spans: 8.0
+19iteration: 15 / gaps: 94 / spans: 4.0
+20iteration: 16 / gaps: 108 / spans: 2.0
+21iteration: 17 / gaps: 117 / spans: 1.0
+22iteration: 18 / gaps: 123 / spans: 0.5
+23iteration: 19 / gaps: 130 / spans: 0.25
+24iteration: 20 / gaps: 134 / spans: 0.125
+25Done with equation #3/785 in 1050 iterations!
+26Saving equation 3 to cache
+Listing 5: Example output from our optimzied input-centric attack tool finding convergence points. Note how after iteration 18 each of the 128 neuron is in it’s own gap.
+Appendix H tlibc’s std::max()) in Tensorflow lite
+1 template <class _T1>
+2 struct __less<_T1, _T1>{
+3 _LIBCPP_INLINE_VISIBILITY _LIBCPP_CONSTEXPR_AFTER_CXX11
+4 bool operator()(const _T1& __x, const _T1& __y) const {return __x < __y;}
+5 };
+6
+7 template <class _Tp, class _Compare>
+8 inline _LIBCPP_INLINE_VISIBILITY _LIBCPP_CONSTEXPR_AFTER_CXX11
+9 const _Tp&
+10 max(const _Tp& __a, const _Tp& __b, _Compare __comp){
+11 return __comp(__a, __b) ? __b : __a;
 12 }
-20 output_data[i] = result;
-21 }
-22}
-Listing 7: Implementation of Logistic() in Tensorflow lite
-Appendix C Error Comparison with Extra Convergence Sets
+Listing 6: std::max() source code from sgxsdk/include/libcxx/algorithm.
+Appendix I std::max() at -Os
+1MOVSS XMM0,dword ptr [param_2]
+2UCOMISS XMM0,dword ptr [param_1]
+3MOV RAX,param_1
+4CMOVA RAX,param_2
+5RET
+Listing 7: Resulting assembly for std::max() at -Os with gcc 11.4.0.
+Appendix J std::max() at -O0
+1PUSH RBP
+2MOV RBP,RSP
+3MOV qword ptr [RBP + local_10],param_1
+4MOV qword ptr [RBP + local_18],param_2
+5MOV RAX,qword ptr [RBP + local_10]
+6MOVSS XMM1,dword ptr [RAX]
+7MOV RAX,qword ptr [RBP + local_18]
+8MOVSS XMM0,dword ptr [RAX]
+9COMISS XMM0,XMM1
+10JBE 001019f0
+11MOV RAX,qword ptr [RBP + local_18]
+12JMP 001019f4
+13MOV RAX,qword ptr [RBP + local_10]
+14POP RBP
+15RET
+Listing 8: Resulting assembly for std::max() at -O0 with gcc 11.4.0.
+Appendix K Error Comparison with Extra Convergence Sets
 Figure 6: Note the effect that extra convergence sets have to reduce variance and lessen the errors, especially after depth=25
+Appendix L Search Strategy Comparison
+Figure 7: A visualisation comparing the execution count of a neuron-centric approach (top) and input-centric approach (bottom) for a single input and 6 normally distributed synthetic ‘neurons’.
+For simplicity, we ignore the calibration process.
+The neuron-centric approach is purely multiplicative (neurons times search depth) so for six neurons and a depth of 10 we cause 60 executions.
+Here, each colour represents one neuron and the lines represent the binary search process.
+The input-centric approach, on the other hand, exploits the distribution of the neurons to ‘save’ executions by recursively bundling queries together for as long as possible until they have diverged into their own unique streams.
+Due to the clustering of these 6 neurons, we can gain the same amount of information in half the number of executions as the first approach.
+In this subplot, colour represents search depth.
+Appendix M Inset of trace from quantised MNIST network
+Figure 8: This is an inset of the region of the trace that corresponds to the 128 neurons in the first layer.
+Note the largely but not entirely constant pattern of steps in the blue-green page at the bottom suggesting possible non-constant time behavior; the first and last values in each "valley" are either 217 or 219 and 77 or 85; everything else appears constant.
+This difference between activations is echoed in the subtly different RIP patterns.